### v1

# Structure of the free interfaces near triple junction singularities in harmonic maps and optimal partition problems

Abstract.
We consider energy-minimizing harmonic maps into trees and we prove the regularity of the singular part of the free interface near triple junction points. Precisely, by a dimension reduction and a topological argument, we prove that there are no admissible frequencies between $1$ and $\nicefrac{{3}}{{2}}$, and we classify the $\nicefrac{{3}}{{2}}$-homogeneous solutions. Then, by proving a new epiperimetric inequality, we show that around any point of frequency $\nicefrac{{3}}{{2}}$, the free interface is composed of three $C^{1,\alpha}$-smooth $(d-1)$-dimensional manifolds (of points of frequency $1$) with common $C^{1,\alpha}$-regular boundary (made of points of frequency $\nicefrac{{3}}{{2}}$) that meet along this boundary at 120 degree angles, which in particular implies that the free interface is a smooth deformation of the $Y$ cone. Our results also apply to spectral optimal partition problems for the Dirichlet eigenvalues. Moreover, as a byproduct, we also prove that the $Y$-configuration is the only minimizer for the min-max optimal $3$-partition problem for the Dirichlet eigenvalues on the sphere, in any dimension.
Key words and phrases:
Triple junction, singular set, epiperimetric inequality, Y-configuration, Almgren’s frequency function, harmonic maps into singular spaces, optimal partitions, segregated systems
1991 Mathematics Subject Classification:
35R35, 49Q10
1. Introduction
This paper is dedicated to the structure of the singular set of harmonic energy-minimizing maps $u:B_{1}\to\Sigma$, defined on $B_{1}\subset\mathbb{R}^{d}$ and with values in singular spaces $\Sigma$. We consider the model case in which the target space is of the form
(1.1) $\Sigma_{N}:=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,,$
By the works of Gromov-Schoen [GS92], Caffarelli-Lin [CL07, CL08], Conti-Terracini-Verzini [CTV02, CTV03, CTV05a] it is known that if
$\mathcal{F}(u)=\{x\in B_{1}\ :\ u(x)=0\},$
is the nodal set of a local minimizer $u=(u_{1},\dots,u_{N})$ of the energy
$E(u,B_{1}):=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x\,,$
then $\mathcal{F}(u)$ can be decomposed into regular and singular parts according to the values of the Almgren’s frequency $\gamma(u,x)$ as follows:
$\mathrm{Reg}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=1\big{\}}\qquad\text{and}\qquad\mathrm{Sing}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>1\big{\}}.$
Gromov-Schoen and Caffarelli-Lin [GS92, CL08] studied the regular part $\mathrm{Reg}(u)$ in every dimension and showed that it is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92]) and is a $(d-2)$-rectifiable ([Alp20, Dee22]). Moreover, Conti, Terracini and Verzini showed in [CTV03] that the singular set in dimension $2$ is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point.
In this paper we study the regularity of the singular set $\mathrm{Sing}(u)$ (and the structure of $\mathcal{F}(u)$ around singular points) in any dimension. Our main result is the following:
Theorem 1.1.
Let $u:B_{1}\to\Sigma_{N}$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$ and $N\geq 1$. Then, the singular set can be decomposed as the disjoint union of
$\mathrm{Sing}(u)=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u):=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, the nodal set $\mathcal{F}(u)$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$.
The main ingredients for the proof of Theorem 1.1 are the classification of the singular homogeneous solutions of lowest frequency, which we achieve in Theorem 1.2 via a topological lemma combined with a dimension reduction argument (see Section 2) and a new epiperimetric inequality (Theorem 1.3) which we obtain via a contradiction argument in the spirit of Weiss [Wei99] (see also [GPS16, FS16]). Furthermore, Theorem 1.1 solves the problem of finding the optimal frequency gap in every dimension $d\geq 2$. Using this result, in Theorem 1.5 we also solve, in any dimension, an open problem by Helffer, Hoffmann-Ostenhof and Terracini in [HHT09], which is the limit case of the Bishop-Friedland-Hayman’s conjecture [Bis92, FH76]. Finally, in Theorem 1.6 and Theorem 1.7, we discuss two corollaries of Theorem 1.1 in the context of harmonic maps and optimal partition problems.
Results in the spirit of Theorem 1.1 exist in the context of area-minimizing currents, where the nodal set $\mathcal{F}(u)$ is replaced by the support of an area-minimizing current (mod $p=3$); in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93], in any dimension $d\geq 3$ (see also the lecture notes [Min24]); we also remark that the analysis for general co-dimension 1 area-minimizing currents mod p was completed only recently in [DHMS20], [DHM+21] and [DHM+22]. To our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$; we also notice that in the framework of Theorem 1.1 the classification of the tangent maps was (until now) an open problem even in dimension $d=3$ and cannot be obtained via elementary geometric arguments. The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, also new in this field.
In the rest of the introduction we give a detailed account on the known results (Section 1.1); in Section 1.2 we state our optimal gap result (Theorem 1.2) and the epiperimetric inequality (Theorem 1.3); finally, Section 1.3 is dedicated to the Helffer-Hoffmann-Ostenhof-Terracini’s problem and the Bishop-Friedland-Hayman’s conjecture.
1.1. Notations and known results
Let us fix the dimension $d\geq 2$. We work with a general open set $D\subseteq\mathbb{R}^{d}$ and with functions $u=(u_{1},\dots,u_{N})\colon D\to\mathbb{R}^{N}$ with $N\geq 2$ components whose support in $D$ is disjoint, i.e. $u_{i}u_{j}\equiv 0$ in $D$. More precisely, we consider the singular (non-positively curved) target space
$\Sigma_{N}:=\Bigg{\{}X\in\mathbb{R}^{N}\colon\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N}X_{i}^{2}X_{j}^{2}=0\Bigg{\}}$ =$=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,,$
on which it is naturally defined the distance
$\mathrm{d}_{\Sigma_{N}}(X,Y):=\begin{cases}|X_{i}-Y_{i}|,&\text{if }X_{j}=Y_{j}=0~{}\text{for all }j\neq i,\\ |X_{i}|+|Y_{j}|,&\text{if }X_{i},Y_{j}>0~{}\text{for some }i\neq j.\end{cases}$
We then consider the Sobolev-type space
$H^{1}(D;\Sigma_{N})=\left\{u\in(H^{1}(D))^{N}\colon u_{i}u_{j}=0~{}\text{a.e. in }D~{}\text{for all }i\neq j\right\}.$
The space $H^{1}_{\textup{loc}}(D;\Sigma_{N})$ is analogously defined. Now, to each open $D\subseteq\mathbb{R}^{d}$ and each function $u\in H^{1}(D;\Sigma_{N})$ we associate the corresponding Dirichlet energy, defined as
$E(u,D):=\sum_{i=1}^{N}\int_{D}|\nabla u_{i}|^{2}\,\mathrm{d}x.$
In the present paper, we work with certain non-negative stationary points of $E$, hence we first define the space of admissible functions as
$\mathcal{A}(D;N):=\left\{u\in H^{1}_{\textup{loc}}(D;\Sigma_{N})\colon u_{i}\geq 0~{}\text{for all }i=1,\dots,N\right\}.$
We consider two classes of “critical points”. The first one is the class of minimizers: we say that $u\in H^{1}_{\textup{loc}}(D;\Sigma_{N})$ is a local minimizer, and we write $u\in\mathcal{M}(D;N)$, if
$E(u,\Omega)\leq E(v,\Omega)$
for all open $\Omega\Subset D$ and all $v\in H^{1}(\Omega;\Sigma_{N})$ such that $u-v\in H^{1}_{0}(\Omega;\Sigma_{N})$. On the other hand, some of our results hold true in a larger class of “critical points”, namely, the so called class $\mathcal{S}$. In order to define it, we first introduce some notation which is also useful in the rest of the paper. Given $u\in\mathcal{A}(D;N)$ and a function
$a\colon\{1,\dots,N\}\to\{-1,1\}$
we define the real-valued function $a\cdot u\in H^{1}_{\textup{loc}}(D)$ as
$(a\cdot u)(x):=\sum_{i=1}^{N}a(i)u_{i}(x)\quad\text{for }x\in D.$
Moreover, for any $i\in\{1,\dots,N\}$, we denote by $\sigma_{i}:=\{1,\dots,N\}\to\{-1,1\}$ the function
$\sigma_{i}(k):=\begin{cases}1,&\text{if }k=i,\\ -1,&\text{if }k\neq 1,\end{cases}$
so that
$\sigma_{i}\cdot u=u_{i}-\sum_{j\neq i}u_{j}\quad\text{in $D$, for all }i=1,\dots,N.$
Now, we say that a function $u\in\mathcal{A}(D;N)$ belongs to the class $\mathcal{S}(D;N)$ if
$-\Delta u_{i}&\leq 0,&&\text{in }D,\\ -\Delta(\sigma_{i}\cdot u)&\geq 0,&&\text{in }D,\end{aligned}\right.$
in the sense of distributions, for all $i=1,\dots,N$. In particular, there holds $\mathcal{M}(D;N)\subseteq\mathcal{S}(D;N)$ (see e.g. [CTV05c, Theorem 5.1]).
The aim of present paper is to investigate regularity questions for solutions $u$ in the classes $\mathcal{S}(D;N)$ and $\mathcal{M}(D;N)$ focusing, in particular, on the geometrical structure and regularity of the free boundary, defined as
$\mathcal{F}(u):=\bigcup_{i=1}^{N}\partial\{x\in D\colon u_{i}(x)>0\},$
in the interior of $D$. We point out that $u\in C^{0,1}_{\textup{loc}}(D;\mathbb{R}^{N})$, see e.g. [CTV05c], which implies that the sets
$\Omega^{u}_{i}:=\{x\in D\colon u_{i}(x)>0\}$
are open and well-defined.
Before stating our main results in Section 1.2, from Section 1.1.1 to Section 1.1.6 we review some known facts concerning critical points $u\in\mathcal{S}(D;N)$. Most of the properties we recall were proved in a series of papers, namely [GS92, Sun03, CL07, Alp20] for the class $\mathcal{M}(D;N)$ and [CTV05c, TT12] for $\mathcal{S}(D;N)$.
1.1.1. Almgren’s frequency function
First of all, for any non-trivial $u\in\mathcal{S}(D;N)$ one can define the Almgren frequency function for any $x_{0}\in D$ and any $r<\operatorname{\mathrm{dist}}(x_{0},\partial D)$ as
$\mathcal{N}(u,r,x_{0}):=\frac{E(u,r,x_{0})}{H(u,r,x_{0})},$
where
$E(u,r,x_{0}):=\frac{1}{r^{d-2}}E(u,B_{r}(x_{0}))=\frac{1}{r^{d-2}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}|\nabla u_{i}|^{2}\,\mathrm{d}x$
is the scaled energy and
$H(u,r,x_{0}):=\frac{1}{r^{d-1}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}u_{i}^{2}\,\mathrm{d}S$
is the scaled height function. The crucial fact is that the $r\mapsto\mathcal{N}(u,r,x_{0})$ is monotone non-decreasing and so the frequency of $u$ at the point $x_{0}$
$\gamma(u,x_{0}):=\lim_{r\to 0^{+}}\mathcal{N}(u,r,x_{0})$
is well-defined. As a consequence, one can derive unique continuation properties implying that the nodal set
$\mathcal{Z}(u)=\{x\in D\colon u_{i}(x)=0~{}\text{for all }i=1,\dots,N\}$
has empty interior and coincides with the free interface, i.e. $\mathcal{Z}(u)=\mathcal{F}(u)$. Moreover, there holds
$\text{either}\quad\gamma(u,x_{0})=1\quad\text{or}\quad\gamma(u,x_{0})\geq 1+\delta_{d},$
for all $x_{0}\in\mathcal{F}(u)$, with $\delta_{d}>0$ being a dimensional constant. Once defined, for any $\gamma\geq 1$,
$\mathcal{F}_{\gamma}(u):=\{x\in\mathcal{F}(u)\colon\gamma(u,x_{0})=\gamma\},$
we split
$\mathcal{F}(u)=\mathrm{Reg}(u)\cup\mathrm{Sing}(u),$
where
$\mathrm{Reg}(u):=\mathcal{F}_{1}(u)\quad\text{and}\quad\mathrm{Sing}(u):=\bigcup_{\gamma\geq 1+\delta_{d}}\mathcal{F}_{\gamma}(u),$
with $\mathrm{Reg}(u)$ relatively open and $\mathrm{Sing}(u)$ relatively closed.
1.1.2. Blow-up analysis
Let $u\in\mathcal{S}(D;N)$, respectively $\mathcal{M}(D;N)$, and let $x_{0}\in\mathcal{F}_{\gamma}(u)$, for some $\gamma\geq 1$. We define the Almgren rescalings as
$u_{x_{0},r}(x):=\frac{u(x_{0}+rx)}{\sqrt{H(u,x_{0},r)}}.$
As a consequence of the Almgren monotonicity formula, by a standard procedure, we know that for any sequence $r_{n}\to 0$ there is a subsequence (still denoted by $r_{n}$) such that
$u_{x_{0},r_{n}}\to U\quad\text{uniformly in $B_{R}$ and in }H^{1}(B_{R};\mathbb{R}^{N})$
as $n\to\infty$, for all $R>0$, for some $U\colon\mathbb{R}^{d}\to\Sigma_{N}$ belonging to
[…]
respectively,
[…]
1.1.3. Points of frequency 1
Let $u\in\mathcal{S}(N,D)$ and let $x_{0}\in\mathcal{F}_{1}(u)$. Then, the limit $U\in\mathcal{S}_{1}(\mathbb{R}^{d};N)$ is unique (i.e. independent from the chosen subsequence) and of the form
$U_{1}=x_{d}^{-},\quad U_{2}=x_{d}^{+},\quad U_{i}\equiv 0~{}\text{for }i=3,\dots,N,$
up to rotations and dilations and there holds
$\frac{u(rx+x_{0})}{r}\to a\,U,~{}\text{uniformly in $B_{R}$ and in $H^{1}(B_{R};\mathbb{R}^{N})$, as }r\to 0,$
for some $a\neq 0$ and all $R>0$. Moreover, around $x_{0}$, the free interface $\mathcal{F}(u)$ is a smooth manifold separating two of the sets $\Omega_{i}^{u}$. Precisely, there is a neighborhood $B_{r}(x_{0})$ and two different indices $i,j\in\{1,\dots,N\}$ such that
$\Omega_{k}^{u}\cap B_{r}(x_{0})=\emptyset\quad\text{for every}\quad k\notin\{i,j\},$
while in $B_{r}(x_{0})$ the function $u_{i}-u_{j}$ is harmonic and
$|\nabla(u_{i}-u_{j})(x_{0})|\neq 0.$
In other words, by choosing $r$ small enough, the interface
$\{u_{i}-u_{j}=0\}\cap B_{r}(x_{0}),$
is smooth $(d-1)$-manifold and we have
$\mathcal{F}(u)\cap B_{r}(x_{0})=\partial\Omega_{i}^{u}\cap B_{r}(x_{0})=\partial\Omega_{j}^{u}\cap B_{r}(x_{0})=\{u_{i}-u_{j}=0\}\cap B_{r}(x_{0}).$
The same statements are true replacing $\mathcal{S}(D;N)$ with $\mathcal{M}(D;N)$ and $\mathcal{S}_{1}(\mathbb{R}^{d};N)$ with $\mathcal{M}_{1}(\mathbb{R}^{d};N)$ (which, actually, coincides with $\mathcal{S}_{1}(\mathbb{R}^{d};N)$).
1.1.4. Classification of blow-ups in dimension 2
In dimension $d=2$, the admissible frequencies and the blow-up limits are completely classified. Namely, it is known that if $x_{0}\in\mathcal{F}(u)$, then
$\gamma(u,x_{0})=\frac{m}{2}\quad\text{for some }m\in\mathbb{N},\ m\geq 2\,.$
Moreover, the sets $\mathcal{S}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ and $\mathcal{M}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ coincide and they contain a unique (up to rotations and dilations) segregated, $\nicefrac{{m}}{{2}}$-homogeneous function. More precisely, if we assume that all the $N$ components are non-trivial, then $N=m$ and any blow-up limit $u\in\mathcal{S}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)=\mathcal{M}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ is of the form
$u_{i}(r,\theta)=\begin{cases}r^{\frac{m}{2}}\left|\cos\left(\frac{m}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{m}(i-1)\leq\theta\leq\pi-\frac{2\pi}{m}i,\\ 0,&\text{elsewhere},\end{cases}$
up to rotations and dilations, for $i=1,\dots,N$.
1.1.5. Clean-up
In [Sun03], the author works in the framework of harmonic maps with $\mathbb{R}$-trees as target space (such as $\Sigma_{N}$). More precisely, in [Sun03, Theorem 1.2] it is proved what, in the language of optimal partition problems, is known as a clean-up result. Once translated in our framework, it says that for any $\gamma\geq 1$ there exists $\varepsilon,\rho>0$ depending on $d$ and $\gamma$ such that if $u\in\mathcal{M}(D;N)$ satisfies
$\mathcal{N}(u,r,x_{0})\leq\gamma\quad\text{and}\quad\left\lVert u_{i}\right\lVert_{L^{\infty}(B_{r}(x_{0}))}\leq\varepsilon\sqrt{H(u,r,x_{0})},$
for some $x_{0}\in D$, $r>0$ (such that $\overline{B_{r}(x_{0})}\subseteq D$) and $i\in I$, with $I\subseteq\{1,\dots,N\}$, then
$u_{i}\equiv 0\quad\text{in }B_{r\rho}(x_{0})\quad\text{for all }i\in I.$
In broad terms, if some components of a minimizer are small in a ball, then they disappear in a smaller ball.
1.1.6. The singular set
By a dimension reduction argument, for any $u\in\mathcal{M}(D;N)$ we have
$\dim_{\mathcal{H}}\mathrm{Sing}(u)\leq d-2.$
Furthermore, thanks to the techniques introduced in [NV17, NV18], in [Alp20, Dee22] it was shown that $\mathrm{Sing}(u)$ is $(d-2)$-rectifiable and that satisfies the following estimate
$|B_{r}(\mathrm{Sing}(u)\cap K)|\leq C(K,u)r^{2},$
for all compact $K\subseteq D$ and all $r>0$ sufficiently small.
1.2. New results on the structure of the singular set
Our main result (Theorem 1.1) provides a fine description of the singular set and of the whole nodal set around points in the lowest stratum. The first question we address is the detection of the optimal frequency gap $\delta_{d}$, defined in Section 1.1.1; until now, the frequency gap constant was explicitly known only in dimension $2$, i.e. $\delta_{2}=\nicefrac{{1}}{{2}}$ (see Section 1.1.4). Hence, our first main result extends this property to any dimension $d\geq 2$. We point out that this is true not only for local minimizers, but for all functions in the class $\mathcal{S}(D;N)$.
Theorem 1.2 (Optimal gap).
Let $u\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ be a non-trivial, $\gamma$-homogeneous critical point, with $\gamma\geq 1$. Then, either
$\gamma=1\quad\text{and}\quad N=2$
or
$\gamma\geq\frac{3}{2}\quad\text{and}\quad N\geq 3.$
Moreover, if $\gamma=\nicefrac{{3}}{{2}}$, then $N=3$ and, up to rotations and dilations, $u$ takes the form of a cylindrical $Y$-configuration, that is
$u(x_{1},\dots,x_{d})=Y(x_{d-1},x_{d}),$
where $Y=(Y_{1},Y_{2},Y_{3})\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{2};3)$ is defined as
$Y_{i}(r,\theta)=\begin{cases}r^{\frac{3}{2}}\left|\cos\left(\frac{3}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{3}(i-1)\leq\theta\leq\pi-\frac{2\pi}{3}i,\\ 0,&\text{elsewhere},\end{cases}$
for $i=1,2,3$.
The above theorem proves that for any $u\in\mathcal{S}(D,N)$ (and, in particular, for any minimizer $u\in\mathcal{M}(D,N)$) the singular set can be decomposed as follows
$\mathrm{Sing}(u)=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
Moreover, every blow-up limit of $u$ at a point $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ is of the form $Y$. This, of course, does not imply that the set
$\mathcal{F}_{\nicefrac{{3}}{{2}}}(u):=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}$
is smooth around $x_{0}$ and not even that the blow-up at $x_{0}$ is unique (as it might depend on the sequence of rescalings $u_{r_{n},x_{0}}$). We next focus on the regularity and the geometrical structure of set $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$. Our approach in tackling this question is through an epiperimetric inequality for $\nicefrac{{3}}{{2}}$-Weiss energy, defined as
$W_{\frac{3}{2}}(u):=E(u,1,0)-\frac{3}{2}H(u,1,0)=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x-\frac{3}{2}\sum_{i=1}^{N}\int_{\partial B_{1}}u_{i}^{2}\,\mathrm{d}S,$
for $u\in H^{1}(B_{1};\Sigma_{N})$. More precisely, we have the following.
Theorem 1.3 (Epiperimetric inequality).
There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on the dimension $d$ such that the following holds. For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that
$\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{Y_{i}>0\}\right)+\sum_{i=4}^{N}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{x_{d-1}=x_{d}=0\}\right)\leq\tau$
and
$\left\lVert\mathrm{d}_{\Sigma_{N}}(c,Y)\right\lVert_{H^{1}(B_{1})}\leq\delta,$
there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u=c$ on $\partial B_{1}$ and
$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
We prove Theorem 1.2 in Section 2. Then, in Section 4 and Section 5 we prove Theorem 1.3. Finally, using this epiperimetric inequality, in Section 6 we prove Theorem 1.1.
1.3. The Bishop-Friedland-Hayman’s conjecture for the min-max
In this section we use Theorem 1.2 to show that the optimal min-max $3$-partition of the sphere is given by the triple junction configuration in any dimension. For any $p\in(0,+\infty]$ we consider the following family of optimal partition problems studied by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10]
($\text{OP}_{p}$) $\min\left\{\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\},$
where $\lambda_{1}(\omega)$ denotes the first Dirichlet eigenvalue of the spherical Laplacian in $\omega\subseteq\partial B_{1}$, and where in the limit case $p=+\infty$, the value
$\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}$
is intended as
$\max_{i\in\{1,\dots,N\}}\lambda_{1}(\omega_{i}).$
In the particular case of the $3$-partition, i.e. $N=3$, we have the following conjecture.
Problem 1.4 (Bishop-Friedland-Hayman-Helffer-Hoffmann-Ostenhof-Terracini).
For $N=3$, prove that the $Y$-configuration is the unique minimizer of ($\text{OP}_{p}$), for any $p\in(0,+\infty]$.
This conjecture was originally posed in [Bis92] in the case $p=1$, but can be generalized for any $p\in(0,+\infty]$ (see [FH76] and [HHT10]). Of particular interest is the case $p=1/2$, which is related with the best exponent in the multi-phase Alt-Caffarelli-Friedman monotonicity formula; in general, there is a vast literature on these optimal partition problems also in relation to models in biology and dynamics of populations, we refer for instance to [CTV03, CTV05b, BBH98, CL07, HHT09, HHT10]. 1.4 was proved for $p=+\infty$ in dimension $d=3$ by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10]; in Section 3 we prove the following theorem, which solves 1.4 (for $p=+\infty$) in any dimension $d\geq 3$.
Theorem 1.5.
The $Y$-configuration is the unique minimizer of
$\min\left\{\max_{i\in\{1,2,3\}}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\}.$
1.4. Two consequences of Theorem 1.1
In this section we discuss two corollaries of Theorem 1.1 for harmonic maps with values in metric graphs and for spectral optimal partitions.
First of all we notice that an immediate corollary of Theorem 1.1 (and the clean-up theorem from [Sun03]), is that the same result holds for harmonic maps with values in a locally finite tree $\Sigma$. Indeed, by [GS92, Sun03, Dee22], if $u:B_{1}\to\Sigma$ is an energy-minimizing map defined in $B_{1}\subset\mathbb{R}^{d}$ with values in a locally finite metric tree, and if $p$ is a vertex of $\Sigma$, then the set $\mathcal{F}:=u^{-1}(\{p\})$ can be decomposed into a regular and a singular parts, $Reg$ and $Sing$, where $Reg$ is a a smooth $(d-1)$-manifold that consists of the points of order (frequency) ${\rm Ord}^{u}$, while $Sing$ has Hausdorff dimension $(d-2)$ and consists of of points of order (frequency) ${\rm Ord}^{u}>1$; furthermore, for every $x\in\mathcal{F}$, there is a neighborhood $\Omega\subset B_{1}$ such that $u:\Omega\to\Sigma$ has values in a subtree $\Sigma_{N}$ of the form (1.1). Thus, we have the following
Theorem 1.6.
Let $\Sigma$ be a metric tree which is locally finite and let $u:B_{1}\to\Sigma$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$. Then, for every vertex $p$ of $\Sigma$ the singular set of $\mathcal{F}:=u^{-1}(\{p\})$ can be decomposed as the disjoint union of
$\mathrm{Sing}=\Big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)>\nicefrac{{3}}{{2}}\Big{\}}.$
The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}:=\big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$, the set $\mathcal{F}$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}$.
Another consequence concerns the regularity of the following optimal partition problem studied in [CL07], [CTV03] and [OV24]. Given a bounded open set $D\subset\mathbb{R}^{d}$, $d\geq 2$, and a natural number $N\geq 2$ we consider the optimal partition problem
(1.2) $\min\Big{\{}\sum_{j=1}^{N}\lambda_{1}(\Omega_{j})\ :\ \Omega_{j}\,-\,\text{open},\ \Omega_{j}\subset D,\ \Omega_{i}\cap\Omega_{j}=\emptyset\ \text{for}\ i\neq j\Big{\}},$
where $\lambda_{1}(\Omega_{j})$ is the first eigenvalue on $\Omega_{j}$ with Dirichlet boundary conditions on $\partial\Omega_{j}$. If the $N$-uple of disjoint sets $(\Omega_{1},\dots,\Omega_{N})$ is a minimizer of (1.2) and if $u=(u_{1},\dots,u_{N})$ is the vector of the normalized first eigenfunctions (each $u_{j}$ is extended by zero outside $\Omega_{j}$), then the free interface
$\mathcal{F}:=D\cap\Big{(}\cup_{j}\partial\Omega_{j}\Big{)}$
can be decomposed as a regular part and a singular part as follows (see [CL07, Alp20]): the regular part $\mathrm{Reg}$ consists of all points $x\in\mathcal{F}$ of frequency $\gamma(u,x)=1$ and is a smooth manifold of dimension $d-1$; the singular part $\mathrm{Sing}$ is a closed $(d-2)$-rectifiable set composed of all the points of frequency $\gamma(u,x)>1$. We have the following theorem
Theorem 1.7.
Let $(\Omega_{1},\dots,\Omega_{N})$ be a minimizer of (1.2) and $u=(u_{1},\dots,u_{N})$ be the corresponding vector of the normalized first eigenfunctions. Then, the singular set of $\mathrm{Sing}$ of the free interface $\mathcal{F}$ can be decomposed as the disjoint union of
$\mathrm{Sing}=\Big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}:=\big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$, the set $\mathcal{F}$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}$.
2. Optimal gap
2.1. A topological lemma
Let $\mathcal{M}$ be a smooth $m$-dimensional manifold. We recall that a subset $\mathcal{F}\subset\mathcal{M}$ is a $C^{1}$ submanifold of dimension $m-1$ if for every $x_{0}\in\mathcal{F}$ there is an open neighborhood $\omega\subset\mathcal{M}$ of $x_{0}$ and a diffeomorphism $\Phi:\omega\to B_{1}\times(-1,1)$ such that $\Phi(\mathcal{F})$ is the graph $\Gamma_{\eta}$ of a $C^{1}$ function $\eta:B_{1}\to(-1,1)$; here $B_{1}$ is the unit ball in $\mathbb{R}^{m-1}$.
Definition 2.1.
Let $\mathcal{M}$ be a smooth $m$-dimensional manifold and let $\mathcal{F}\subset\mathcal{M}$ be a $C^{1}$ submanifold of dimension $m-1$. Given $x\in\mathcal{M}\setminus\mathcal{F}$ and $y\in\mathcal{M}\setminus\mathcal{F}$, we denote by $C(x\to y)$ the family of piecewise $C^{1}$ curves transversal to $\mathcal{F}$. Precisely, we say that $\ell\in\mathcal{C}(x\to y)$ if:
  1. (1)
    $\ell\colon[0,1]\to\mathcal{M}$, $\ell(0)=x$ and $\ell(1)=y$;
  2. (2)
    $\ell\in C([0,1];\mathcal{M})$ and there are $k=k(\ell)$ points $0=T_{0}<T_{1}<\dots<T_{k}=1,$ such that: $\ell\in C^{1}((T_{j-1},T_{j});\mathcal{M})$ and $\ell^{\prime}\neq 0$ on $(T_{j-1},T_{j})$ for every $j=1,\dots,k$, and
    $\ell(T_{j})\notin\mathcal{F}\quad\text{for every}\quad j={0},\dots,k;$
  3. (3)
    if $\ell(t)\in\mathcal{F}$ for some $t\in(0,1)$, then $\ell$ is transversal to $\mathcal{F}$ at $\ell(t)$, that is, $\ell^{\prime}(t)\notin T_{\ell(t)}\mathcal{F}$, where $T_{\ell(t)}\mathcal{F}$ is the tangent space to $\mathcal{F}$ at $\ell(t)$.
Lemma 2.2.
Let $\mathcal{M}$ be a smooth manifold of dimension $m$. Suppose that $\mathcal{F}$ is an $(m-1)$-dimensional $C^{1}$ submanifold of $\mathcal{M}$ and suppose that $\mathcal{F}$ is a relatively closed subset of $\mathcal{M}$. Let $x,y\in\mathcal{M}\setminus\mathcal{F}$ and let $\ell\in\mathcal{C}(x\to y)$ be a curve in $\mathcal{M}$. Then, the set of intersection points ():=$\mathcal{I}(\ell):=\{t\in(0,1)\ :\ \ell(t)\in\mathcal{F}\}$ is finite.
Proof.
Suppose that the set $\mathcal{I}$ contains an infinite sequence $(t_{n})_{n\geq 1}$. Then, we can suppose that $t_{n}$ converges to some $t_{\infty}\in(0,1)$. By continuity of $\gamma$ and the fact that $\mathcal{F}$ is closed, we have $\gamma(t_{\infty})\in\mathcal{I}$. But $\gamma$ is transversal to $\mathcal{F}$ at $t_{\infty}$. Thus, for $t\neq t_{\infty}$ in a neighborhood of $t_{\infty}$, $\gamma(t)\notin\mathcal{F}$, which is a contradiction. ∎
Lemma 2.3.
Let $\mathcal{M}$ be a simply connected smooth manifold of dimension $m$. Suppose that $\mathcal{F}\subset\mathcal{M}$ is a relatively closed subset of $\mathcal{M}$ and a $C^{1}$-submanifold of dimension $m-1$. Suppose that $x\in\mathcal{M}\setminus\mathcal{F}$ and that $\ell:[0,1]\to\mathcal{M}$ is a closed curve in $\mathcal{C}(x\to x)$. Then the set $\mathcal{I}(\ell)$ has an even number of elements.
Proof.
Since $\mathcal{M}$ is simply connected, we can find a continuous homotopy $\Gamma:[0,1]\times[0,1]\to\mathcal{M}$ that deforms $\ell$ into the constant curve $x$. Precisely,
$\Gamma(0,\cdot)=\ell(\cdot)\qquad\text{and}\qquad\Gamma(1,\cdot)\equiv x.$
We proceed in several steps.
Step 1. Discrete homotopy. Since $\mathcal{F}$ is a $C^{1}$ embedded submanifold, we can find points $(s_{i},t_{j})$ of the form
$(s_{i},t_{j})=\left(\frac{i}{n},\frac{j}{n}\right)\quad i=0,\dots,n\ ,\ j=1,\dots,n,$
such that for every set of “consecutive” 4 points:
$Q_{i,j}:=\Big{\{}X_{i,j}:=\Gamma(s_{i},t_{j}),~{}X_{i+1,j}:=\Gamma(s_{i+1},t_{j}),~{}X_{i,j+1}:=\Gamma(s_{i},t_{j+1}),~{}X_{i+1,j+1}:=\Gamma(s_{i+1},t_{j+1})\Big{\}}$
we have that either:
  1. (Q1)
    $Q_{i,j}$ is contained in an open set $\omega_{1}\subset\mathcal{M}\setminus\mathcal{F}$ diffeomorphic to a ball;
  2. (Q2)
    $Q_{i,j}$ is contained in an open set $\omega_{2}\subset\mathcal{M}$ for which there is a diffeomorphism $\Phi:\omega_{2}\to B_{1}\times(-1,1)$ such that $\Phi(\mathcal{F})$ is the graph $\Gamma_{\eta}$ of a $C^{1}$ function $\eta:B_{1}\to(-1,1)$, $B_{1}$ being the unit ball in $\mathbb{R}^{m-1}$.
We next notice that by perturbing slightly each of the points $X_{i,j}$ we can construct another family of $n\times n$ points
[…]
such that any set of the form
$\widetilde{Q}_{i,j}:=\Big{\{}\widetilde{X}_{i,j},\ \widetilde{X}_{i+1,j},\ \widetilde{X}_{i,j+1},\ \widetilde{X}_{i+1,j+1}\Big{\}}\,,$
satisfies (Q1) or (Q2) above. Moreover, we can also suppose that
$\widetilde{X}(0,j)\in\ell([0,1])\quad\text{for every}\quad j=0,\dots,n;$
$\widetilde{X}(n,j)=x\quad\text{for every}\quad j=0,\dots,n.$
Step 2. Construction of a piecewise $C^{1}$-regular net with vertices $\widetilde{X}_{ij}$. In what follows we will use the following notation, for every $0\leq i,j\leq n$
  • $\omega_{ij}$ is the open set
    $\omega_{ij}:=\begin{cases}\omega_{1}\quad\text{if}\ (Q1)\ \text{is verified for the set}\ \widetilde{Q}_{ij}\,;\\ \omega_{2}\quad\text{if}\ (Q2)\ \text{is verified for the set}\ \widetilde{Q}_{ij}\,.\end{cases}$
  • $\alpha_{ij}\in\mathcal{C}(\widetilde{X}_{i,j}\to\widetilde{X}_{i,j+1})$ is a curve lying in $\omega_{ij}$ (so $\alpha_{ij}$ does not cross $\mathcal{F}$);
  • $\beta_{ij}^{+}\in\mathcal{C}(\widetilde{X}_{i,j}\to\widetilde{X}_{i+1,j})$ is a curve lying in $\omega_{ij}$ and $\beta_{ij}^{-}(t):=\beta_{ij}^{+}(1-t)$.
Finally, for every $i=1,\dots,n$, we set
$\ell_{i}:=\alpha_{i1}\ast\alpha_{i2}\ast\dots\ast\alpha_{i(n-1)},$
where $\ast$ denotes the usual concatenation of curves. We notice that, up to a reparametrization, we can choose the curves $\alpha_{0j}$, $j=1,\dots,n$, to be pieces of $\ell$. Moreover, we can choose $\alpha_{nj}$, $j=1,\dots,n$, to be the constant curves $\alpha_{nj}(t)=x$ for all $t$. Thus, for all $t$, we have:
$\ell_{0}(t)=\ell(t)\qquad\text{and}\qquad\ell_{n}(t)=x.$
Step 3. The map $M$. In what follows, for any curve $\ell:[0,1]\to\mathcal{M}$ with $\ell\in\mathcal{C}(x\to y)$ for some $x,y\in\mathcal{M}\setminus\mathcal{F}$, we will denote by $M(\ell)$ the number of elements of $\mathcal{I}(\ell)$. The map $M$ has the following properties:
  1. (M1)
    Let $x,y,z\in\mathcal{M}\setminus\mathcal{F}$. If $\ell_{1}\in\mathcal{C}(x\to y)$ and $\ell_{2}\in\mathcal{C}(y\to z)$, then $\ell_{1}\ast\ell_{2}\in\mathcal{C}(x\to z)$ and
    $M(\ell_{1}\ast\ell_{2})=M(\ell_{1})+M(\ell_{2}).$
  2. (M2)
    Let $x,y\in\mathcal{M}\setminus\mathcal{F}$. If $\ell_{+}\in\mathcal{C}(x\to y)$ and $\ell_{-}$ is defined as
    $\ell_{-}(t)=\ell(1-t),$
    then $\ell_{-}\in\mathcal{C}(y\to x)$ and $M(\ell_{+})=M(\ell_{-})$.
  3. (M3)
    Let $x,y,z,w\in\mathcal{M}\setminus\mathcal{F}$ be 4 points in a set $\omega_{2}$ as in (Q2) and let the curves $\alpha,\beta,\gamma,\delta$ be as follows:
    $\alpha\in\mathcal{C}(x\to y)\ ,\quad\beta\in\mathcal{C}(y\to z)\ ,\quad\gamma\in\mathcal{C}(z\to w)\ ,\quad\delta\in\mathcal{C}(x\to w)\ .$
    Then, we have that
    $M(\delta)=M(\alpha\ast\beta\ast\gamma)=M(\alpha)+M(\beta)+M(\gamma).$
Step 4. Conclusion. Now, using the properties (M1), (M2) and (M3), we obtain that, for every pair of consecutive indices $i$ and $i+1$ with $0\leq i<i+1\leq n$, the following holds:
$M(\ell_{i+1})=\sum_{j=0}^{n-1}M(\alpha_{(i+1)j})$ $=\sum_{j=0}^{n-1}M(\beta_{ij}^{-}\ast\alpha_{ij}\ast\beta_{ij}^{+})$
$=\sum_{j=0}^{n-1}\Big{(}M(\beta_{ij}^{-})+M(\alpha_{ij})+M(\beta_{ij}^{+})\Big{)}$
$=\sum_{j=0}^{n-1}\Big{(}M(\alpha_{ij})+2M(\beta_{ij}^{+})\Big{)}$
$=M(\ell_{i})+2\sum_{j=0}^{n-1}M(\beta_{ij}^{+}),$
so that
$(-1)^{M(\ell_{i+1})}=(-1)^{M(\ell_{i})}.$
Since $i$ is arbitrary, this implies
$(-1)^{M(\ell_{n})}=(-1)^{M(\ell_{0})}=(-1)^{M(\ell)}.$
Finally, since $\ell_{n}$ is the constant curve $\ell_{n}(t)=x$ for every $t$, we get that $M(\ell_{n})=0$. This concludes the proof since $M(\ell)$ by definition is the number of elements of $\mathcal{I}(\ell)$. ∎
2.2. Proof of Theorem 1.2
In this section we prove Theorem 1.2. The proof is based on a dimension reduction argument and on the following
Lemma 2.4.
Let $d\geq 3$ and $N\geq 2$. Suppose that $u=(u_{1},\dots,u_{N})\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ is a non-trivial $\gamma$-homogeneous solution of the optimal partition problem, i.e.
$u(x)=|x|^{\gamma}u\left(\frac{x}{|x|}\right)\quad\text{and}\quad\left\{\begin{aligned} -\Delta u_{i}&\leq 0,&&\text{in }\mathbb{R}^{d},\\ -\Delta(\sigma_{i}\cdot u)&\geq 0,&&\text{in }\mathbb{R}^{d},\end{aligned}\right.$
in a distributional sense, for all $i=1,\dots,N$. Suppose that all the points on the unit sphere are regular, that is
$\mathcal{F}(u)\equiv\mathcal{F}_{1}(u)\quad\text{on}\quad\partial B_{1},$
and that all the sets $\Omega_{i}^{u}\cap\partial B_{1}$, $i=1,\dots,N$ are connected. Then, there is a function
$\sigma:\{1,\dots,N\}\to\{-1,1\},$
such that the function $\sigma\cdot u$, given by
$(\sigma\cdot u)(x):=\sum_{i=1}^{N}\sigma(i)u_{i}(x)$
is harmonic in $\mathbb{R}^{d}$.
Proof.
Without loss of generality we can suppose that
$\Omega_{i}^{u}\cap\partial B_{1}\neq\emptyset\quad\text{for every}\quad i=1,\dots,N.$
We define the function $\sigma$ as follows. We set $\sigma(1)=1$ and we fix a point $x_{0}\in\Omega_{1}^{u}\cap B_{r}$. For every point $x_{1}\in\partial B_{1}\setminus\mathcal{F}(u)$ we take a curve $\ell\in\mathcal{C}(x_{0}\to x_{1})$ (where $\mathcal{C}(x_{0}\to x_{1})$ is as in Definition 2.1 with $\mathcal{M}=\partial B_{1}$ and $\mathcal{F}=\mathcal{F}(u)$) and we define
$\sigma(x_{1}):=(-1)^{M(\ell)},$
$M(\ell)$ being the number of times $\ell$ crosses $\mathcal{F}(u)$:
(2.1) $M(\ell):=\sharp\Big{\{}t\in[0,1]\ :\ \ell(t)\in\mathcal{F}(u)\Big{\}}.$
By Lemma 2.3 (applied to the sphere $\mathcal{M}=\partial B_{1}$ and $\mathcal{F}=\mathcal{F}(u)$) the value of $\sigma(x_{1})$ does not depend on the choice of the curve $\ell$. In order to prove the harmonicity of $\sigma\cdot u$ we notice that if $B_{r}(x_{0})$ is a ball such that $\mathcal{F}(u)\cap B_{r}(x_{0})=\partial\Omega_{i}\cap\partial\Omega_{j}\cap B_{r}(x_{0})$ is a smooth graph in $B_{r}(x_{0})$, then for every pair of points $x_{i}\in\Omega_{i}\cap B_{r}(x_{0})$ and $x_{j}\in\Omega_{j}\cap B_{r}(x_{0})$, we can find a curve in $\mathcal{C}(x_{i}\to x_{j})$ that crosses $\mathcal{F}(u)$ only once. Thus, by construction
$\sigma(x_{i})=-\sigma(x_{j})$
and so the function $\sigma\cdot u$ is harmonic across $\mathcal{F}(u)$. This concludes the proof of Lemma 2.4. ∎
We are now in position to prove our first main theorem.
Proof of Theorem 1.2.
In any dimension $d\geq 2$ and for any $N\geq 2$, we define:
(2.2) $A(d,N):=\inf\Big{\{}\gamma>1\ :\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)\neq\emptyset\Big{\}}.$
Since the admissible frequencies in dimension $d=2$ are classified, we have:
$A(2,N)=\frac{3}{2}\qquad\text{and}\qquad 1<A(d,N)\leq\frac{3}{2}\quad\text{when}\quad d\geq 2.$
We will show that $A(d,N)=3/2$ for every $d\geq 2$ by proving that
(2.3) $A(d-1,N)\leq A(d,N)\quad\text{for all }d\geq 2.$
In order to do this, we proceed in three steps.
Step 1. The infimum in (2.2) is reached. Indeed, suppose that there is a sequence of $\gamma_{n}$-homogeneous global solutions $u^{n}\in\mathcal{S}_{\gamma_{n}}(\mathbb{R}^{d};N)$ such that $\gamma_{n}\to A(d,N)$. Up to a rescaling, we suppose that:
$\|u^{n}\|_{L^{2}(\partial B_{1})}=1\qquad\text{and}\qquad\|\nabla u^{n}\|_{L^{2}(B_{1})}=\gamma_{n}.$
Up to extracting a subsequence, we can suppose that $u^{n}$ converges strongly in $L^{2}(B_{1};\mathbb{R}^{N})$ and $L^{2}(\partial B_{1};\mathbb{R}^{N})$, and weakly in $H^{1}(B_{1};\mathbb{R}^{N})$ to some $A(d,N)$-homogeneous global solution $u\in\mathcal{S}_{A(d,N)}(\mathbb{R}^{d},N)$.
Step 2. $\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}\neq\emptyset.$ Take any $k\in\{1,\dots,N\}$ and any connected component $\omega_{k}$ of $\Omega_{k}$. Since the component $u_{k}$ is homogeneous and non-negative in $\mathbb{R}^{d}$, its trace on $\omega_{k}\cap\partial B_{1}$ is (a multiple of) the first eigenfunction of the spherical Laplacian on $\omega_{k}\cap\partial B_{1}$ with eigenvalue
$\lambda:=\gamma(\gamma+d-2)\quad\text{where}\quad\gamma=A(d,N),$
with Dirichlet boundary conditions on $\partial\omega_{k}\cap\partial B_{1}$. In particular, by the Faber-Krahn inequality, the area of each $\omega_{k}\cap\partial B_{1}$ is bounded from below by a constant depending only on $\gamma$ and $d$, so the number of connected components of $\Omega_{k}$ is finite. Let $\widetilde{N}$ be the number of all connected components of $\mathbb{R}^{d}\setminus\mathcal{F}(u)$. We define the $\gamma$-homogeneous $\widetilde{u}\in\mathcal{A}(\mathbb{R}^{d};\widetilde{N})$ as the vector whose components are the restrictions of $u$ to the connected components of $\Omega_{k}$, $k=1,\dots,N$. It is immediate to check that $\widetilde{u}$ is still a global solution of the optimal partition problem, i.e. $\widetilde{u}\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};\widetilde{N})$. We now claim that $\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}\neq\emptyset$. If not, since by construction all $\widetilde{\Omega}_{j}:=\{\widetilde{u}_{j}>0\}$, $1\leq j\leq\widetilde{N}$ are connected, we can apply Lemma 2.4; thus, there is a map $\sigma:\{1,\dots,\widetilde{N}\}\to\{-1,1\}$ such that $\sigma\cdot\widetilde{u}$ is harmonic and $A(d,N)$-homogeneous in $\mathbb{R}^{d}$. Thus $A(d,N)$ is an integer and, since $A(d,N)>1$, we get that $A(d,N)\geq 2$. This is a contradiction with the bound $A(d,N)\leq 3/2$.
Step 3. Dimension reduction. From the previous step we know that there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$. Without loss of generality, we can assume that $x_{0}=\vec{\bm{e}}_{1}$. By the definition of $A(d,N)$ we know that
$A(d,N)\leq\gamma(u,\vec{\bm{e}}_{1}).$
On the other hand, the homogeneity of $u$ implies that
$\gamma(u,t\vec{\bm{e}}_{1})=\gamma(u,\vec{\bm{e}}_{1})\quad\text{for every}\quad t>0.$
In particular, by the upper semicontinuity of $\gamma(u,\cdot)$ and the homogeneity of $u$, we have
$A(d,N)=\gamma(u,0)\geq\lim_{t\to 0^{+}}\gamma(u,t\vec{\bm{e}}_{1})=\gamma(u,\vec{\bm{e}}_{1}).$
Thus, we get
$A(d,N)=\gamma(u,0)=\gamma(u,\vec{\bm{e}}_{1}),$
but this implies that $u$ is invariant in the direction of $\vec{\bm{e}}_{1}$, that is:
$u(x+t\vec{\bm{e}}_{1})=u(x)\quad\text{for every}\quad t\in\mathbb{R}.$
But then the nontrivial function
$\widetilde{u}(x_{2},\dots,x_{d}):=u(0,x_{2},\dots,x_{d}),$
belongs to $\mathcal{S}_{A(d;N)}(\mathbb{R}^{d-1};N)$. This implies that $A(d-1;N)\leq A(d;N)$, thus proving (2.3) and so (2.2).
In order to conclude the proof of Theorem 1.2, we are left to classify the $\nicefrac{{3}}{{2}}$-homogeneous global solutions. Let $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$. By Step 3 there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$. But then, reasoning as in Step 3, we have that
$\gamma(u,0)=\gamma(u,x_{0})=\frac{3}{2}$
and that $u$ is invariant in the direction of $x_{0}$. Up to a rotation of the coordinate system, we can suppose that $x_{0}=\vec{\bm{e}}_{1}$. Then $\widetilde{u}(x_{2},\dots,x_{d}):=u(0,x_{2},\dots,x_{d})$ is a $\nicefrac{{3}}{{2}}$-homogeneous global solution in $\mathbb{R}^{d-1}$, i.e. it belongs to $\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d-1};N)$. By repeating this argument in every dimension $d-k$, $k\geq 1$, we finally get that, up to a rotation of the coordinate axes, we have
$u(x_{1},\dots,x_{d-1},x_{d})=w(x_{d-1},x_{d}),$
where $w$ is a $\nicefrac{{3}}{{2}}$-homogeneous solution in dimension 2. This gives the specific form of $w$ and the fact that $w$ has exactly three non-zero components. ∎
3. Optimal partition of the sphere
The purpose of this section is to prove Theorem 1.5, essentially basing ourselves on the results obtained in Section 2, namely Theorem 1.2 (and in [CTV05b, HHT09]). By [CTV05b, Theorem 3.1] we know that there exists $\phi\in H^{1}(\partial B_{1};\Sigma_{3})\cap C^{0,1}(B_{1};\mathbb{R}^{3})$ such that $\phi_{i}\not\equiv 0$ and $(\omega_{1},\omega_{2},\omega_{3})$, defined as $\omega_{i}:=\{\phi_{i}>0\}$, achieve
(3.1) $\mathcal{L}_{3}:=\min\left\{\max_{i=1,2,3}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\}.$
Moreover, [CTV05b, Theorem 3.1] it is also proved that $\phi\in\mathcal{S}(\partial B_{1},3)$, in the sense that
$-\Delta_{\partial B_{1}}\phi_{i}&\leq 0,&&\text{in }\partial B_{1},\\ -\Delta_{\partial B_{1}}(\sigma_{i}\cdot\phi)&\geq 0,&&\text{in }\partial B_{1}.\end{aligned}\right.$
Now we recall [HHT09, Remark 3.14] which says that all the first eigenvalues are equal, namely $\lambda_{1}(\omega_{i})=\mathcal{L}_{3}$ for all $i=1,2,3$; we then let $\gamma>0$ be such that
$\gamma(\gamma+d-2)=\mathcal{L}_{3}.$
We now consider the $\gamma$-homogeneous extensions of $\phi_{i}$, that is $\varphi_{i}(r,\theta):=r^{\gamma}\phi_{i}(\theta)$. One can easily check that $\varphi\in H^{1}_{\textup{loc}}(\mathbb{R}^{d};\Sigma_{3})\cap C^{0,1}_{\textup{loc}}(\mathbb{R}^{d};\mathbb{R}^{3})$ belongs to the class $\mathcal{S}_{\gamma}(\mathbb{R}^{d},3)$, i.e.
$-\Delta\varphi_{i}&\leq 0,&&\text{in }\mathbb{R}^{d},\\ -\Delta(\sigma_{i}\cdot\varphi)&\geq 0,&&\text{in }\mathbb{R}^{d}.\end{aligned}\right.$
At this point, the proof of Theorem 1.5 follows easily. Indeed, one can test (3.1) with the $Y$-configuration and directly obtain that
$\mathcal{L}_{3}\leq\frac{3}{2}\left(\frac{3}{2}+d-2\right),$
thus implying that $\gamma\leq\nicefrac{{3}}{{2}}$. On the other hand, since $N=3$, by applying Theorem 1.2 to $\varphi$, we obtain that $\gamma=\nicefrac{{3}}{{2}}$ and that $\varphi$ coincides with the $Y$-configuration.
4. Linearization around triple junctions
We now introduce the linearized problem at points of frequency $\nicefrac{{3}}{{2}}$, which plays a crucial role in the proof of the epiperimetric inequality. For $i=1,2,3$, we set
$\Omega_{i}:=\left\{(r\cos\theta,r\sin\theta)\colon r>0~{}\text{and }\theta\in\left(\pi-(i-1)\frac{2\pi}{3},\pi-i\frac{2\pi}{3}\right)\right\},$
so that $\mathbb{R}^{2}$ is the disjoint union
$\mathbb{R}^{2}=\Omega_{1}\cup\Omega_{2}\cup\Omega_{3}\cup Y.$
Moreover, we may identify $\Omega_{i}$ with its cylindrical extension $\mathbb{R}^{d-2}\times\Omega_{i}$. Now, the linearized problem around the triple junction solution $Y=(Y_{1},Y_{2},Y_{3})\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{2};3)$, which we recall to be defined as
$Y_{i}(r,\theta)=\begin{cases}r^{\frac{3}{2}}\left|\cos\left(\frac{3}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{3}(i-1)\leq\theta\leq\pi-\frac{2\pi}{3}i,\\ 0,&\text{elsewhere},\end{cases}$
is the following:
(4.1) $-\Delta w_{k}&=0,&&\text{in }\Omega_{k}~{}\text{for every }k\in\{1,2,3\},\\ w_{i}&=-w_{j},&&\text{on }\partial\Omega_{i}\cap\partial\Omega_{j},~{}\text{for every }i\neq j\in\{1,2,3\},\\ \nabla w_{i}&=-\nabla w_{j}&&\text{on }\partial\Omega_{i}\cap\partial\Omega_{j},~{}\text{for every }i\neq j\in\{1,2,3\}.\end{aligned}\right.$
More precisely, we say that $w=(w_{1},w_{2},w_{3})$ is a solution of (4.1) if
$\begin{cases}w_{i}-w_{j}\in H^{1}_{\textup{loc}}(\Omega_{ij}),\\ -\Delta(w_{i}-w_{h})=0\quad\text{in }\Omega_{ij},\end{cases}$
where
$\Omega_{ij}:=\Omega_{i}\cup\Omega_{j}\cup\left(\partial\Omega_{i}\cap\partial\Omega_{j}\right)\setminus\{0\}$
for all $i,j\in\{1,2,3\}$, $i\neq j$.
We remark that the linearized problem can also be rephrased with the cut space as domain, drawing a parallel with the thin obstacle problem. Namely, if we let
$\mathcal{P}:=\{(x^{\prime\prime},x_{d-1},x_{d})\in\mathbb{R}^{d}\colon x_{d-1}\leq 0,~{}x_{d}=0\}$
and $\widehat{w}:=-w_{1}+w_{2}-w_{3}$, then we have the following:
  • $\widehat{w}\in H^{1}_{\textup{loc}}(\mathbb{R}^{d}\setminus\mathcal{P})$ solves
    $-\Delta w&=0,&&\text{in }\mathbb{R}^{d}\setminus\mathcal{P},\\ \lim_{x_{d}\to 0^{+}}w+\lim_{x_{d}\to 0^{-}}w&=0,&&\text{on }\mathcal{P},\\ \lim_{x_{d}\to 0^{+}}\partial_{x_{d}}w+\lim_{x_{d}\to 0^{-}}\partial_{x_{d}}w&=0,&&\text{on }\mathcal{P};\end{aligned}\right.$
  • we can write $\widehat{w}=w_{\textup{e}}+w_{\textup{o}}$, where $w_{\textup{e}}$ and $w_{\textup{o}}$ are, respectively, even and odd with respect to $x_{d}$:
    $w_{\textup{e}}(x^{\prime},x_{d}):=\frac{w(x^{\prime},x_{d})+w(x^{\prime},-x_{d})}{2},\qquad w_{\textup{o}}(x^{\prime},x_{d}):=\frac{w(x^{\prime},x_{d})-w(x^{\prime},-x_{d})}{2};$
  • the even part $w_{\textup{e}}$ solves
    $-\Delta w_{\textup{e}}&=0,&&\text{in }\mathbb{R}^{d}\setminus\mathcal{P},\\ w_{\textup{e}}&=0,&&\text{on }\mathcal{P};\end{aligned}\right.$
  • the odd part $w_{\textup{o}}$ solves
    $-\Delta w_{\textup{o}}&=0,&&\text{in }\mathbb{R}^{d}\setminus\mathcal{P},\\ \partial_{x_{d}}w_{\textup{o}}&=0,&&\text{on }\mathcal{P}.\end{aligned}\right.$
A key observation is that, if we consider the even extension of the restriction of $w_{\textup{o}}$ to +$\mathbb{R}^{d}_{+}$ and exchange $x_{d-1}$ with $-x_{d-1}$, that is we let
$\widetilde{w}(x^{\prime\prime},x_{d-1},x_{d}):=\begin{cases}w_{\textup{o}}(x^{\prime\prime},-x_{d-1},x_{d}),&\text{for }x_{d}>0,\\ w_{\textup{o}}(x^{\prime\prime},-x_{d-1},-x_{d}),&\text{for }x_{d}\leq 0,\end{cases}$
then one can see that $\widetilde{w}$ solves the same equation as the even part $w_{\textup{e}}$, namely
$-\Delta\widetilde{w}&=0,&&\text{in }\mathbb{R}^{d}\setminus\mathcal{P},\\ \widetilde{w}&=0,&&\text{on }\mathcal{P}.\end{aligned}\right.$
On the other hand, this coincides with the linearized problem of the thin obstacle problem and so, for what concerns the classification of $\nicefrac{{3}}{{2}}$-homogeneous solutions, we can appeal to the known results in this setting (see e.g. [GPS16, FS16]). Therefore, we proved the following.
Proposition 4.1.
Let $\widehat{w}\in H^{1}_{\textup{loc}}(\mathbb{R}^{d}\setminus\mathcal{P})$ be a $\nicefrac{{3}}{{2}}$-homogeneous solution of
$-\Delta w&=0,&&\text{in }\mathbb{R}^{d}\setminus\mathcal{P},\\ \lim_{x_{d}\to 0^{+}}w+\lim_{x_{d}\to 0^{-}}w&=0,&&\text{on }\mathcal{P},\\ \lim_{x_{d}\to 0^{+}}\partial_{x_{d}}w+\lim_{x_{d}\to 0^{-}}\partial_{x_{d}}w&=0,&&\text{on }\mathcal{P}.\end{aligned}\right.$
Then $\widehat{w}=w_{\textup{e}}+w_{\textup{o}}$, where $w_{\textup{e}}$ is even with respect to $x_{d}$ and takes the form
$w_{\textup{e}}=a\widehat{Y}(x_{d-1},x_{d})+U_{0}(x_{d-1},x_{d})\sum_{i=1}^{d-2}a_{i}x_{i}$
and $w_{\textup{o}}$ is odd with respect to $x_{d}$ and takes the form
$w_{\textup{o}}=b\widehat{Y}(-x_{d-1},x_{d})+U_{0}(-x_{d-1},x_{d})\sum_{i=1}^{d-2}b_{i}x_{i}\quad\text{for }x_{d}>0,$
where
$\widehat{Y}(r,\theta)=r^{\frac{3}{2}}\cos\left(\frac{3}{2}\theta\right)=-Y_{1}+Y_{2}-Y_{3}\quad\text{and}\quad U_{0}(r,\theta)=r^{\frac{1}{2}}\cos\left(\frac{\theta}{2}\right).$
In particular, $w_{\textup{e}}$ and $w_{\textup{o}}$ are $\nicefrac{{3}}{{2}}$-homogeneous solutions of
$-\Delta w_{\textup{e}}&=0,&&\text{in }\mathbb{R}^{d}\setminus\mathcal{P},\\ w_{\textup{e}}&=0,&&\text{on }\mathcal{P}\end{aligned}\right.\qquad\text{and}\qquad\left\{\begin{aligned} -\Delta w_{\textup{o}}&=0,&&\text{in }\mathbb{R}^{d}\setminus\mathcal{P},\\ \partial_{x_{d}}w_{\textup{o}}&=0,&&\text{on }\mathcal{P}.\end{aligned}\right.$
5. Epiperimetric inequality at points of frequency $\nicefrac{{3}}{{2}}$
Lemma 5.1 (Distance on $\Sigma_{N}$).
For any $u,v\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u_{i},v_{i}\geq 0$ in $B_{1}$ for all $i=1,\dots,N$, we have
$\mathrm{d}_{\Sigma_{N}}(u,v)=\Big{|}v_{i}-u_{i}+\sum_{j\neq i}u_{j}\Big{|}=\left|v_{i}-\sigma_{i}\cdot u\right|\quad\text{in }\overline{\{v_{i}>0\}},$
Moreover, we also have
$\mathrm{d}_{\Sigma_{N}}(u,v)=\sum_{j=1}^{N}|u_{j}-v_{j}|\quad\text{in }B_{1}.$
Proof.
Let $x\in\overline{\{v_{i}>0\}}$. If $x\in\{u_{i}>0\}$, then
$\mathrm{d}_{\Sigma_{N}}(u,v)(x)=|v_{i}(x)-u_{i}(x)|=\Big{|}v_{i}(x)-u_{i}(x)+\sum_{j\neq i}u_{j}\Big{|}=\sum_{j=1}^{N}|v_{j}(x)-u_{j}(x)|,$
while, if $x\in\overline{\{u_{k}>0\}}$, with $k\neq i$, then
$\mathrm{d}_{\Sigma_{N}}(u,v)(x)=v_{i}+u_{k}=\Big{|}v_{i}(x)-u_{i}(x)+\sum_{j\neq i}u_{j}(x)\Big{|}=\sum_{j=1}^{N}|v_{j}(x)-u_{j}(x)|.$
If $u_{k}(x)=0$ for all $k=1,\dots,N$ the statement is trivial. ∎
Lemma 5.2.
Suppose that $v\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$. Then, for every $u\in C^{0,1}(B_{1})$ and every $i=1,\dots,N$, we have that
$\int_{B_{1}\cap\{v_{i}>0\}}\nabla v_{i}\cdot\nabla u\,\mathrm{d}x=\gamma\int_{\partial B_{1}\cap\{v_{i}>0\}}v_{i}u\,\mathrm{d}S+\int_{B_{1}\cap\partial\{v_{i}>0\}}u\partial_{\bm{\nu}}v_{i}\,\mathrm{d}S,$
where $\nu$ is the normal to $\mathrm{Reg}(v)\cap\partial\{v_{i}>0\}$ pointing outwards $\{v_{i}>0\}$.
Proof.
Before starting, we first recall that we know some basic structure of the free boundary of $v$; more precisely, from [TT12] we know that $v\in C^{0,1}(B_{1};\mathbb{R}^{N})$ and that we can decompose
$\mathcal{F}(v)=\bigcup_{i=1}^{N}\partial\{v_{i}>0\}=\mathrm{Reg}(v)\cup\mathrm{Sing}(v),$
where $\mathrm{Reg}(v)$ is the union of $(d-1)$-dimensional manifolds of class $C^{1,\alpha}$ and $\mathrm{Sing}(v)$ has zero $\mathcal{H}^{d-1}$ measure. Moreover,
(5.1) $|\nabla v_{i}|=|\nabla v_{j}|\quad\text{on }\mathrm{Reg}(v)\qquad\text{and}\quad|\nabla v_{i}|=0\quad\text{on }\mathrm{Sing}(v).$
Hence, for every $\varepsilon>0$ we can find a finite family of balls $B_{r_{k}}(x_{k})$ such that:
${\rm Sing}(v)\cap\overline{B}_{1}\subset\bigcup_{k}B_{r_{k}}(x_{k})\qquad\text{and}\qquad\sum_{k}r_{k}^{d-1}\leq\varepsilon.$
For every ball we consider a smooth function $\phi_{k}\in C^{\infty}_{c}(\mathbb{R}^{d})$ such that:
$\phi_{k}\equiv 1\ \text{ on }\ B_{r_{k}}(x_{k}),\qquad\phi_{k}\equiv 0\ \text{on }\ \mathbb{R}^{d}\setminus B_{2r_{k}}(x_{k}),\qquad|\nabla\phi_{k}|\leq 2r_{k}^{-1}.$
and we set $\phi=\phi_{\varepsilon}:=\sup_{k}\phi_{k}$. Then, since $u(1-\phi)$ is identically zero in a neighborhood of $\mathrm{Sing}(v)$, we have that
$\int_{B_{1}\cap\{v_{i}>0\}}\nabla\big{(}u(1-\phi)\big{)}\cdot\nabla v_{i}=\gamma\int_{\partial B_{1}\cap\{v_{i}>0\}}(1-\phi)v_{i}u\,\mathrm{d}S+\int_{B_{1}\cap\partial\{v_{i}>0\}}(1-\phi)u\partial_{\bm{\nu}}v_{i}\,\mathrm{d}S.$
Thus, in order to conclude the proof, it is sufficient to estimate the error term
$e:=\int_{B_{1}\cap\{v_{i}>0\}}(\phi\nabla u+u\nabla\phi)\cdot\nabla v_{i}+\gamma\int_{\partial B_{1}\cap\{v_{i}>0\}}\phi v_{i}u\,\mathrm{d}S+\int_{B_{1}\cap\partial\{v_{i}>0\}}\phi u\partial_{\bm{\nu}}v_{i}\,\mathrm{d}S.$
By the definition of $\phi$ we have that
$|e|$ $\leq\sum_{k}\Big{(}\int_{B_{2r_{k}}(x_{k})\cap B_{1}\cap\{v_{i}>0\}}|\phi_{k}||\nabla u||\nabla v_{i}|+\int_{B_{2r_{k}}(x_{k})\cap B_{1}\cap\{v_{i}>0\}}|\nabla\phi_{k}||u||\nabla v_{i}|$
$\qquad+\gamma\int_{B_{2r_{k}}(x_{k})\cap\partial B_{1}\cap\{v_{i}>0\}}|\phi_{k}||v_{i}||u|\,\mathrm{d}S+\int_{B_{2r_{k}}(x_{k})\cap B_{1}\cap\partial\{v_{i}>0\}}|\phi_{k}||u||\nabla v_{i}|\,\mathrm{d}S\Big{)}$
$\leq C\sum_{k}\Big{(}r_{k}^{d}+r_{k}^{d}r_{k}^{-1}+\gamma r_{k}^{d-1}+r_{k}^{d-1}\Big{)}\leq C\sum_{k}r_{k}^{d-1}\leq C\varepsilon,$
where the constant $C$ depends on $d$, $\|v\|_{L^{\infty}}$, $\|\nabla v\|_{L^{\infty}}$, $\|u\|_{L^{\infty}}$, and $\|\nabla u\|_{L^{\infty}}$. ∎
Lemma 5.3.
For any $u\in H^{1}(B_{1};\Sigma_{N})$ and any $\gamma\geq 0$, we let
$W_{\gamma}(u):=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x-\gamma\sum_{i=1}^{N}\int_{\partial B_{1}}u_{i}^{2}\,\mathrm{d}S.$
Then, for any $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ and $v\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ such that $u_{i},v_{i}\geq 0$ for all $i=1,\dots,N$, there holds
$W_{\gamma}(u)=W_{\gamma}(\mathrm{d}_{\Sigma_{N}}(u,v))+\beta(u,v),$
where
(5.2) $\beta(u,v):=4\sum_{1\leq i<j\leq N}\int_{B_{1}\cap\partial\{v_{i}>0\}\cap\partial\{v_{j}>0\}}|\nabla v_{i}|\Big{(}\sum_{k\neq i,j}u_{k}\Big{)}\,\mathrm{d}S.$
Proof.
We prove the result by direct computations. In view of Lemma 5.1 we have that
$W_{\gamma}(\mathrm{d}_{\Sigma_{N}}(u,v))$ $=\sum_{i=1}^{N}\Bigg{[}\int_{B_{1}\cap\{v_{i}>0\}}\Big{|}\nabla\Big{(}v_{i}-u_{i}+\sum_{k\neq i}u_{k}\Big{)}\Big{|}^{2}\,\mathrm{d}x$
$\quad-\gamma\int_{\partial B_{1}\cap\{v_{i}>0\}}\Big{|}v_{i}-u_{i}+\sum_{k\neq i}u_{k}\Big{|}^{2}\,\mathrm{d}S\Bigg{]}$
and so, by explicit computations, we derive that
$W_{\gamma}(\mathrm{d}_{\Sigma_{N}}(u,v))$ $=W_{\gamma}(u)+W_{\gamma}(v)$
$-2\sum_{i=1}^{N}\int_{B_{1}\cap\{v_{i}>0\}}\nabla v_{i}\cdot\nabla\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}x+2\gamma\int_{\partial B_{1}\cap\{v_{i}>0\}}v_{i}\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}S.$
Now, in view of Lemma 5.2, (5.1) and by homogeneity, we have that
$\sum_{i=1}^{N}\int_{B_{1}\cap\{v_{i}>0\}}\nabla v_{i}\cdot\nabla\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}x$ $=\sum_{i=1}^{N}\int_{B_{1}\cap\partial\{v_{i}>0\}}\partial_{\bm{\nu}}v_{i}\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}x$
$\quad+\sum_{i=1}^{N}\int_{\partial B_{1}}\partial_{\bm{\nu}}v_{i}\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}S$
$=-\sum_{i=1}^{N}\int_{B_{1}\cap\partial\{v_{i}>0\}}|\nabla v_{i}|\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}x$
$\quad+\gamma\sum_{i=1}^{N}\int_{\partial B_{1}}v_{i}\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}S$
Therefore, since also $W_{\gamma}(v)=0$ (again from Lemma 5.2), we have
$W_{\gamma}(\mathrm{d}_{\Sigma_{N}}(u,v))$ $=W_{\gamma}(u)+2\sum_{i=1}^{N}\int_{B_{1}\cap\partial\{v_{i}>0\}}|\nabla v_{i}|\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}x$
$=W_{\gamma}(u)+2\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\int_{B_{1}\cap\partial\{v_{i}>0\}\cap\partial\{v_{j}>0\}}|\nabla v_{i}|\Big{(}u_{i}-\sum_{k\neq i}u_{k}\Big{)}\,\mathrm{d}x$
$=W_{\gamma}(u)-4\sum_{1\leq i<j\leq N}\int_{B_{1}\cap\partial\{v_{i}>0\}\cap\partial\{v_{j}>0\}}|\nabla v_{i}|\Big{(}\sum_{k\neq i,j}u_{k}\Big{)}\,\mathrm{d}x,$
thus concluding the proof. ∎
Theorem 5.4.
There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on $d$ such that the following holds. For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $c_{n,i}\geq 0$ for all $i=1,\dots,N$,
$\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{Y_{i}>0\}\right)+\sum_{i=4}^{N}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{x_{d-1}=x_{d}=0\}\right)\leq\tau$
and
$\left\lVert\mathrm{d}_{\Sigma_{N}}(c,Y)\right\lVert_{H^{1}(B_{1})}\leq\delta,$
there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u_{i}\geq 0$ for all $i$, $u=c$ on $\partial B_{1}$, and
$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
Proof.
In the whole proof, with an abuse of notation, for the sake of simplicity, we write $\Omega_{i}$ in place of $\Omega_{i}\cap B_{1}$, for $i=1,2,3$.
We reason by contradiction, and assume that there exist sequences $\{\delta_{n}\}_{n}\subseteq\mathbb{R}_{+}$, $\{\varepsilon_{n}\}_{n}\subseteq\mathbb{R}_{+}$, $\{\tau_{n}\}_{n}\subseteq\mathbb{R}_{+}$, $\{N_{n}\}_{n}\subseteq\mathbb{N}$ and $\{c_{n}\}_{n}\subseteq H^{1}(B_{1};\Sigma_{N_{n}})\cap C^{0,1}(B_{1};\mathbb{R}^{N_{n}})$ such that
$c_{n,i}\geq 0\quad\text{for all }n\in\mathbb{N}~{}\text{and all }i=1,\dots,N_{n},$
$\delta_{n},\varepsilon_{n},\tau_{n}\to 0\quad\text{as }n\to\infty,$
(5.3) $\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\left(\{c_{n,i}>0\},\{Y_{i}>0\}\right)+\sum_{i=4}^{N_{n}}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{x_{d-1}=x_{d}=0\}\right)=\tau_{n},$
$\left\lVert\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},Y)\right\lVert_{H^{1}(B_{1})}=\delta_{n}$
and
(5.4) $(1-\varepsilon_{n})W_{\frac{3}{2}}(c_{n})<W_{\frac{3}{2}}(u)$
for all $u\in H^{1}(B_{1};\Sigma_{N_{n}})\cap C^{0,1}(B_{1};\mathbb{R}^{N_{n}})$ such that $u=c_{n}$ on $\partial B_{1}$. Moreover, without loss of generality, we can assume that
$\delta_{n}=\left\lVert\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},Y)\right\lVert_{H^{1}(B_{1})}=\inf\left\{\left\lVert\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},h))\right\lVert_{H^{1}(B_{1})}\colon h\in\mathcal{M}_{\frac{3}{2}}(\mathbb{R}^{d};3)\right\}.$
Now, first of all, we apply Lemma 5.3 with $v=Y$ and divide both sides by $\delta_{n}^{2}$, thus obtaining that
$(1-\varepsilon_{n})\left(W_{\frac{3}{2}}(\xi_{n})+\frac{1}{\delta_{n}^{2}}\beta(c_{n},Y)\right)<W_{\frac{3}{2}}\left(\frac{\mathrm{d}_{\Sigma_{N_{n}}}(u,Y)}{\delta_{n}}\right)+\frac{1}{\delta_{n}^{2}}\beta(u,Y),$
for all $u\in H^{1}(B_{1};\Sigma_{N_{n}})\cap C^{0,1}(B_{1};\mathbb{R}^{N_{n}})$ such that $u=c_{n}$ on $\partial B_{1}$, where $\xi_{n}:=\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},Y)/\delta_{n}$. In other words,
$(1-\varepsilon_{n})\left(\int_{B_{1}}|\nabla\xi_{n}|^{2}\,\mathrm{d}x+\frac{1}{\delta_{n}^{2}}\beta(c_{n},Y)\right)<\int_{B_{1}}\left|\nabla\Big{(}\frac{\mathrm{d}_{\Sigma_{N_{n}}}(u,Y)}{\delta_{n}}\Big{)}\right|^{2}\,\mathrm{d}x+\frac{1}{\delta_{n}^{2}}\beta(u,Y)+\frac{3\varepsilon_{n}}{2}\int_{\partial B_{1}}\xi_{n}^{2}\,\mathrm{d}S.$
We also let
$w_{n,i}:=\begin{cases}\displaystyle\frac{Y_{i}-c_{n,i}}{\delta_{n}}+\sum_{j\neq i}\frac{c_{n,j}}{\delta_{n}},&\text{in }\Omega_{i},\\ 0,&\text{in }B_{1}\setminus\Omega_{i},\end{cases}$
for $i=1,2,3$. Since
$\left|w_{n,i}\right|=\frac{\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},Y)}{\delta_{n}}=\xi_{n},\quad\text{in }\Omega_{i},$
see Lemma 5.1, in particular,
$\sum_{i=1}^{3}\left\lVert w_{n,i}\right\lVert_{H^{1}(\Omega_{i})}^{2}=\left\lVert\xi_{n}\right\lVert_{H^{1}(B_{1})}^{2}=1.$
Therefore, there exist functions $w_{i}\in H^{1}(\Omega_{i})$ such that, up to a subsequence,
$w_{n,i}\rightharpoonup w_{i}\quad\text{weakly in }H^{1}(\Omega_{i}),~{}\text{as }n\to\infty,$
for $i=1,2,3$. Hence, we can rephrase the contradiction assumption as
(5.5) $(1-\varepsilon_{n})\left(\sum_{i=1}^{3}\int_{\Omega_{i}}|\nabla w_{n,i}|^{2}\,\mathrm{d}x+\frac{1}{\delta_{n}^{2}}\beta(c_{n},Y)\right)\\ <\int_{B_{1}}\Big{|}\nabla\Big{(}\frac{\mathrm{d}_{\Sigma_{N_{n}}}(u,Y)}{\delta_{n}}\Big{)}\Big{|}^{2}\,\mathrm{d}x+\frac{1}{\delta_{n}^{2}}\beta(u,Y)+o(1),$
as $n\to\infty$, for all $u\in H^{1}(B_{1};\Sigma_{N_{n}})\cap C^{0,1}(B_{1};\mathbb{R}^{N_{n}})$ such that $u=c_{n}$ on $\partial B_{1}$, where the reminder term $o(1)$ is independent from $u$. We now proceed in the following steps, which, at the end, lead to a contradiction:
  • Step 1: $\delta_{n}^{-2}\beta(c_{n},Y)$ is uniformly bounded with respect to $n$;
  • Step 2: $w=(w_{1},w_{2},w_{3})$ is a solution of the linearized problem, see Section 4;
  • Step 3: $w_{i}=0$ for all $i=1,2,3$;
  • Step 4: $w_{n,i}\to 0$ strongly in $H^{1}(\Omega_{i})$, as $n\to\infty$, for all $i=1,2,3$.
Proof of Step 1. For any $n\in\mathbb{N}$, we define a suitable “competitor” $u_{n}\in H^{1}(B_{1};\Sigma_{N_{n}})$ such that $u_{n}=c_{n}$ on $\partial B_{1}$ as follows. First, in an inner ball, we let $u_{n}=Y$, namely
$u_{n}=Y\quad\text{in }B_{\frac{1}{2}}.$
In the outer annulus $B_{1}\setminus B_{\frac{1}{2}}$, we consider the interpolation between $Y$ and $c_{n}$; namely, we first set
$t=t_{r}:=2\left(r-\frac{1}{2}\right),\quad\text{with }r=|x|.$
Now, we let
$u_{n}(r,\theta)=r^{\frac{3}{2}}\left(t\,c_{n,i}(1,\theta)+(1-t)Y_{i}(1,\theta)\right)\,\vec{\bm{e}}_{i}$
if $11/2<r<$ 1 and $Y_{j}(\theta)=c_{n,j}(1,\theta)=0$ for all $j\neq i$ and, on the other hand, we let
$u_{n}(r,\theta)=\begin{cases}r^{\frac{3}{2}}(Y_{i}(1,\theta)-t(Y_{i}(1,\theta)+c_{n,j}(1,\theta)))\,\vec{\bm{e}}_{i},&\text{for }0<t<\frac{Y_{i}(1,\theta)}{Y_{i}(1,\theta)+c_{n,j}(1,\theta)},\\ r^{\frac{3}{2}}(-Y_{i}(1,\theta)+t(Y_{i}(1,\theta)+c_{n,j}(1,\theta)))\,\vec{\bm{e}}_{j},&\text{for }\frac{Y_{i}(1,\theta)}{Y_{i}(1,\theta)+c_{n,j}(1,\theta)}\leq t<1,\end{cases}$
if $11/2<r<$ 1 and $Y_{i}(\theta)>0$ and $c_{n,j}(1,\theta)>0$ for some $j\neq i$. If we now let
$\widetilde{u}_{n,i}:=\frac{Y_{i}-u_{n,i}+\sum_{j\neq i}u_{n,j}}{\delta_{n}}\quad\text{in }\Omega_{i},$
then there holds
$\left|\widetilde{u}_{n,i}\right|=\frac{\mathrm{d}_{\Sigma_{N_{n}}}(u_{n},Y)}{\delta_{n}}\quad\text{in }\Omega_{i},$
see Lemma 5.1, and one can easily check that
$\widetilde{u}_{n,i}=tw_{n,i}\quad\text{in }\Omega_{i}\cap\left(B_{1}\setminus B_{\frac{1}{2}}\right)\quad\text{and}\quad\widetilde{u}_{n,i}=0\quad\text{in }B_{\frac{1}{2}}.$
Moreover, we have that
$u_{n}=tc_{n}\quad\text{on }\{(Y_{1},Y_{2},Y_{3})=(0,0,0)\}.$
Therefore, from (5.5) we derive that
(5.6) $\frac{\beta((1-\varepsilon_{n}-t)c_{n},Y)}{\delta_{n}^{2}}<(\varepsilon_{n}-1)\sum_{i=1}^{3}\int_{\Omega_{i}}|\nabla w_{n,i}|^{2}\,\mathrm{d}x+\sum_{i=1}^{3}\int_{\Omega_{i}}|\nabla(tw_{n,i})|^{2}\,\mathrm{d}x+o(1),$
as $n\to\infty$. On one hand, being $w_{n,i}$ bounded in $H^{1}(\Omega_{i})$, the right hand side is uniformly bounded with respect to $n\in\mathbb{N}$, while, on the other hand, by direct computations we have
(5.7) $\beta((1-\varepsilon_{n}-t)c_{n},Y)=(d+1)\int_{0}^{1}r^{d}(1-\varepsilon_{n}-t)\,\mathrm{d}r\,\beta(c_{n},Y).$
Therefore since $t<1$ in $B_{1}\setminus B_{1/2}$, from (5.6) we conclude the proof of the step.
Proof of Step 2. We are going to prove that
(5.8) $-\Delta(w_{i}-w_{j})=0\quad\text{in }\Omega_{ij},$
for all $i,j=1,2,3$, $i\neq j$, where
$\Omega_{ij}:=\left(\Omega_{i}\cup\Omega_{j}\cup\left(\partial\Omega_{i}\cap\partial\Omega_{j}\right)\right)\cap B_{1}.$
For the sake of simplicity, we prove it for $i=1$ and $j=2$. First of all, we fix any $x_{0}\in\Omega_{12}$ and $r>0$ in such a way that $\overline{B_{2r}(x_{0})}\subseteq\Omega_{12}$. Moreover, for fixed $\tau>0$, we let
$C_{\tau}:=\left\{(x^{\prime\prime},x_{d-1},x_{d})\in\mathbb{R}^{d}\colon\sqrt{x_{d-1}^{2}+x_{d}^{2}}<\tau\right\}$
and we assume that $B_{2r}(x_{0})\cap C_{\tau}=\emptyset$ and that $c_{n,j}=0$ in $B_{2r}(x_{0})$, for $j=3,\dots,N_{n}$. This is possible by choosing $\tau$ sufficiently small and $n$ sufficiently large, in view of the Hausdorff convergence assumption (5.3). We now use (5.5) with $u\in H^{1}(B_{1};\Sigma_{N_{n}})$ being such that $u=c_{n}$ in $B_{1}\setminus B_{2r}(x_{0})$ and $u_{j}=0$ in $B_{2r}(x_{0})$ for $j=3,\dots,N_{n}$. Since $c_{n,j}=u_{n,j}\equiv 0$ in $B_{2r}(x_{0})$, we have that
$\beta(u,Y)=\beta(c_{n},Y),$
which, in view of Step 1, gives
(5.9) $\frac{1}{\delta_{n}^{2}}\beta(u,Y)-\frac{1-\varepsilon_{n}}{\delta_{n}^{2}}\beta(c_{n},Y)=\frac{\varepsilon_{n}}{\delta_{n}^{2}}\beta(c_{n},Y)=o(1),$
as $n\to+\infty$. Moreover, we observe that
$\frac{\mathrm{d}_{\Sigma_{N_{n}}}(u_{n},Y)}{\delta_{n}}=\frac{1}{\delta_{n}}\Bigg{|}Y_{i}-c_{n,i}+\sum_{j\neq i}c_{n,j}\Bigg{|}=w_{n,i}\quad\text{in }\Omega_{i}\setminus B_{2r}(x_{0}),$
for $i=1,2,3$. Now, from (5.5) and (5.9) we deduce that
$\int_{\Omega_{1}\cap B_{2r}(x_{0})}|\nabla w_{n,1}|^{2}\,\mathrm{d}x+\int_{\Omega_{2}\cap B_{2r}(x_{0})}|\nabla w_{n,2}|^{2}\,\mathrm{d}x+o(1)\\ <\int_{\Omega_{1}\cap B_{2r}(x_{0})}\Bigg{|}\nabla\left(\frac{Y_{1}-u_{1}+u_{2}}{\delta_{n}}\right)\Bigg{|}^{2}\,\mathrm{d}x+\int_{\Omega_{2}\cap B_{2r}(x_{0})}\Bigg{|}\nabla\left(\frac{Y_{2}-u_{2}+u_{1}}{\delta_{n}}\right)\Bigg{|}^{2}\,\mathrm{d}x,$
for all $u\in H^{1}(B_{1};\Sigma_{2})$ such that $u=c_{n}$ on $\partial B_{2r}(x_{0})$, where the reminder term does not depend on $u$. In particular, if we let
$W_{n}:=w_{n,1}-w_{n,2}\quad\text{and}\quad U:=u_{1}-u_{2}$
this implies that
(5.10) $\int_{B_{2r}(x_{0})}|\nabla W_{n}|^{2}\,\mathrm{d}x<\int_{B_{2r}(x_{0})}\Bigg{|}\nabla\Bigg{(}\frac{Y_{1}-Y_{2}-U}{\delta_{n}}\Bigg{)}\Bigg{|}^{2}\,\mathrm{d}x+o(1)$
for all $U\in H^{1}(B_{2r}(x_{0}))$ such that $U=c_{n,1}-c_{n,2}$ on $\partial B_{2r}(x_{0})$. One can also see that, by definition,
$w_{n,1}=-w_{n,2}\quad\text{and}\quad\nabla w_{n,1}=-\nabla w_{n,2}\quad\text{on }\left(\partial\Omega_{1}\cap\partial\Omega_{2}\right)\cap B_{2r}(x_{0}),$
so that $W_{n}\in H^{1}(B_{2r}(x_{0}))\cap C^{0,1}(B_{2r}(x_{0}))$; moreover, its $H^{1}$ norm is bounded and so, up to a subsequence,
$W_{n}\rightharpoonup w_{1}-w_{2}\quad\text{weakly in }H^{1}(B_{2r}(x_{0})),~{}\text{as }n\to\infty.$
At this point, we define
$U=\eta(Y_{1}-Y_{2}-\delta_{n}\phi)+(1-\eta)(c_{n,1}-c_{n,2}),$
where $\eta\in C_{c}^{\infty}(B_{2r}(x_{0}))$ is such that $\eta=1$ in $B_{r}(x_{0})$ and $\phi\in H^{1}(B_{2r}(x_{0}))$. Plugging this choice into (5.10), we get
$\int_{B_{2r}(x_{0})}|\nabla W_{n}|^{2}\,\mathrm{d}x<\int_{B_{2r}(x_{0})}|\nabla((1-\eta)W_{n}+\eta\phi)|^{2}\,\mathrm{d}x+o(1).$
By direct computations, one can see that
$\int_{B_{2r}(x_{0})}(1-(1-\eta)^{2})|\nabla W_{n}|^{2}\,\mathrm{d}x+o(1)\\ <\int_{B_{2r}(x_{0})}\left(W_{n}^{2}|\nabla\eta|^{2}-2(1-\eta)W_{n}\nabla W_{n}\cdot\nabla\eta+|\nabla(\eta\phi)|^{2}+2\nabla((1-\eta)W_{n})\cdot\nabla(\eta\phi)\right)\,\mathrm{d}x$
and, since we can pass now to the limit as $n\to\infty$ in the inequality above, after rearranging the terms back we obtain that
$\int_{B_{2r}(x_{0})}|\nabla(w_{1}-w_{2})|^{2}\,\mathrm{d}x\leq\int_{B_{2r}(x_{0})}|\nabla((1-\eta)(w_{1}-w_{2})+\eta\phi)|^{2}\,\mathrm{d}x.$
Finally, we take $\phi=w_{1}-w_{2}$ in $B_{2r}(x_{0})\setminus B_{r}(x_{0})$, so that
$\int_{B_{r}(x_{0})}|\nabla(w_{1}-w_{2})|^{2}\,\mathrm{d}x\leq\int_{B_{r}(x_{0})}|\nabla\phi|^{2}\,\mathrm{d}x.$
Hence, we proved that $\Delta(w_{1}-w_{2})=0$ in $B_{r}(x_{0})$. Up to moving the ball, we get (5.8).
Proof of Step 3. We use the classification of the $\nicefrac{{3}}{{2}}$-homogeneous solutions of the linearized problem, namely Proposition 4.1. More precisely, if we denote
$\widehat{w}:=-w_{1}+w_{2}-w_{3},$
which, we recall, belongs to $H^{1}(B_{1}\setminus\mathcal{P})$, then in view of Proposition 4.1 we have that
$\widehat{w}=w_{\textup{e}}+w_{\textup{o}},$
where $w_{\textup{e}}$ is even with respect to $x_{d}$ and is of the form
$w_{\textup{e}}(x)=a_{0}\widehat{Y}(x_{d-1},x_{d})+U_{0}(x_{d-1},x_{d})\sum_{j=1}^{d-2}a_{j}x_{j},\quad\text{for }x_{d}>0,$
for some $a_{j}\in\mathbb{R}$, $j=0,\dots,d-2$, and $w_{\textup{o}}$ is odd with respect to $x_{d}$ and is of the form
$w_{\textup{o}}(x)=b_{0}\widehat{Y}(-x_{d-1},x_{d})+U_{0}(-x_{d-1},x_{d})\sum_{j=1}^{d-2}b_{j}x_{j}\quad\text{for }x_{d}>0,$
for some $b_{j}\in\mathbb{R}$, $j=0,\dots,d-2$, where
$\widehat{Y}(r,\theta)=r^{\frac{3}{2}}\cos\left(\frac{3}{2}\theta\right)=-Y_{1}+Y_{2}-Y_{3}\quad\text{and}\quad U_{0}(r,\theta)=r^{\frac{1}{2}}\cos\left(\frac{\theta}{2}\right).$
Moreover, if we denote by $Z$ and $V_{0}$, respectively, the odd extensions of $\widehat{Y}(-x_{d-1},x_{d})$ and $U_{0}(-x_{d-1},x_{d})$ and if we pass to polar coordinates in the last two variables, i.e.
{$x_{d-1}&=r\cos\theta\\ x_{d}&=r\sin\theta,\end{aligned}\right.\qquad\text{with }r\in(0,1)~{}\text{and}~{}\theta\in\left(-\pi,\pi\right),$
then we have
(5.11) $Z(r,\theta)=-r^{\frac{3}{2}}\sin\left(\frac{3}{2}\theta\right)\quad\text{and}\quad V_{0}(r,\theta)=r^{\frac{1}{2}}\sin\left(\frac{\theta}{2}\right).$
We also observe that, by explicit computations
(5.12) $\partial_{x}\widehat{Y}(r,\theta)=\frac{3}{2}r^{\frac{1}{2}}\cos\left(\frac{\theta}{2}\right)\quad\text{and}\quad\partial_{y}\widehat{Y}(r,\theta)=-\frac{3}{2}r^{\frac{1}{2}}\sin\left(\frac{\theta}{2}\right).$
We now prove that all the coefficients $a_{j},b_{j}$, with $j=0,\dots,d-2$ vanish, by exploiting the fact that $Y$ is the projection of $c_{n}$ onto the space of blow-up limits $\mathcal{M}_{\frac{3}{2}}(\mathbb{R}^{d};3)$. We know that
$\left\lVert\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},Y)\right\lVert_{H^{1}(B_{1})}\leq\left\lVert\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},h)\right\lVert_{H^{1}(B_{1})}\quad\text{for all }h\in\mathcal{M}_{\frac{3}{2}}(\mathbb{R}^{d};3)$
and, since
$\mathrm{d}_{\Sigma_{N_{n}}}(c_{n},h)=\Big{|}h_{i}-c_{n,i}+\sum_{j\neq i}c_{n,j}\Big{|}\quad\text{on }\Omega_{i},$
we can rephrase it as follows
(5.13) $\sum_{i=1}^{3}\left(w_{n,i},\frac{Y_{i}-h_{i}}{\delta_{n}}\right)_{H^{1}(\Omega_{i})}\leq\frac{\delta_{n}}{2}\sum_{i=1}^{3}\left\lVert\frac{h_{i}-Y_{i}}{\delta_{n}}\right\lVert_{H^{1}(\Omega_{i})}^{2}\quad\text{for all }h\in\mathcal{M}_{\frac{3}{2}}(\mathbb{R}^{d};3).$
At this point, we consider as $h$ suitable perturbations of the blow-up $Y$ (depending on the parameter $\delta_{n}$) and pass to the limit as $n\to\infty$. This will provide a set of orthogonality conditions which the limit $w$ must satisfy; in view of the classification, this leads to conditions on the coefficients $a_{j}$ and $b_{j}$. More precisely, we are going to choose a multiple of the blow-up limit $Y$ and the infinitesimal rotations inside the coordinate planes $x_{i}x_{j}$, with $i=1,\dots,d-1$ and $j=d-1,d$ ($i\neq j$). This way, we get a total of $2(d-2)$ equations which coincide with the degrees of freedom of the problem. We proceed in the following steps.
  1. (1)
    $a_{0}=0$: we choose $h(x)=(1\pm\delta_{n})Y(x_{d-1},x_{d})$.
  2. (2)
    $a_{j}=0$ for $j=1,\dots,d-2$: we choose $h(x)=Y(x_{d-1}\cos(\delta_{n})\pm x_{i}\sin(\delta_{n}),x_{d})$.
  3. (3)
    $b_{0}=0$: we choose $h(x)=Y(x_{d-1}\cos(\delta_{n})\mp x_{d}\sin(\delta_{n}),x_{d}\cos(\delta_{n})\pm x_{d-1}\sin(\delta_{n}))$.
  4. (4)
    $b_{j}=0$ for $j=1,\dots,d-2$: we choose $h(x)=Y(x_{d-1},x_{d}\cos(\delta_{n})\pm x_{j}\sin(\delta_{n}))$.
We observe that perturbations (1) and (2) leave the last coordinate $x_{d}$ untouched, and so the generate conditions only on the even part of $w$. Let us now perform the various computations.
  1. (1)
    Plugging $h$ into (5.13), we get that
    $\mp\sum_{i=1}^{3}\left(w_{n,i},Y_{i}\right)_{H^{1}(\Omega_{i})}\leq\frac{\delta_{n}}{2}\sum_{i=1}^{3}\left\lVert Y_{i}\right\lVert_{H^{1}(\Omega_{i})}^{2},$
    and, passing to the limit as $n\to\infty$, this leads to
    $\sum_{i=1}^{3}(\widehat{w},\widehat{Y})_{H^{1}(\Omega_{i})}=(\widehat{w},\widehat{Y})_{H^{1}(B_{1}\setminus\mathcal{P})}=0,$
    being $\widehat{w},\widehat{Y}\in H^{1}(B_{1}\setminus\mathcal{P})$. In view of the classification of the solutions of the linearized problem, since $\widehat{Y}$ is even with respect to $x_{d}$ and independent from $x_{j}$, for $j=1,\dots,d-2$, the terms in $\widehat{w}$ which linearly depend on $x_{j}$ do not contribute; thus, we have
    $a_{0}\|\widehat{Y}\|_{H^{1}(B_{1}\setminus\mathcal{P})}^{2}=(\widehat{w},\widehat{Y})_{H^{1}(B_{1}\setminus\mathcal{P})}=0,$
    which implies that $a_{0}=0$.
  2. (2)
    First of all, we observe that
    $\frac{\widehat{Y}(x_{d-1}\cos(\delta_{n})\pm x_{j}\sin(\delta_{n}),x_{d})-\widehat{Y}(x_{d-1},x_{d})}{\delta_{n}}\to\pm\frac{3}{2}x_{j}U_{0}(x_{d-1},x_{d})$
    strongly in $H^{1}(\Omega_{i})$, as $n\to\infty$, for all $i=1,2,3$ and $j=1,\dots,d-2$. Therefore, plugging $h$ into (5.13) and passing to the limit as $n\to\infty$ yields
    $\sum_{i=1}^{3}(\widehat{w},x_{j}U_{0})_{H^{1}(\Omega_{i})}=(\widehat{w},x_{j}U_{0})_{H^{1}(B_{1}\setminus\mathcal{P})}=0.$
    On the other hand, since $\widehat{Y}$ is even with respect to $x_{d}$ and since $x_{j}U_{0}$ is independent from the variable $x_{k}$ for $k=1,\dots,d-2$, $k\neq j$, thanks to the explicit form of $\widehat{w}$ we infer
    $(\widehat{w},x_{j}U_{0})_{H^{1}(B_{1}\setminus\mathcal{P})}=a_{j}\left\lVert x_{j}U_{0}\right\lVert_{H^{1}(B_{1}\setminus\mathcal{P})}^{2}=0,$
    thus proving that $a_{j}=0$.
  3. (3)
    By (5.11) and (5.12) we deduce that
    $\frac{\widehat{Y}(x_{d-1}\cos(\delta_{n})\mp x_{d}\sin(\delta_{n}),x_{d}\cos(\delta_{n})\pm x_{d-1}\sin(\delta_{n}))-\widehat{Y}(x_{d-1},x_{d})}{\delta_{n}}$
    converges to ±$\pm Z(x_{d-1},x_{d})$ weakly in $H^{1}(\Omega_{i})$, for $i=1,2,3$, as $n\to\infty$. Hence, reasoning analogously to the previous step, from (5.13) we get
    $\left(\widehat{w},Z\right)_{H^{1}(B_{1}\setminus\mathcal{P})}=0.$
    Since the second term of the scalar product is odd with respect to $x_{d}$ and since the right term does not depend on $x_{j}$ for any $j=1,\dots,d-2$, in view also of the explicit form of $\widehat{w}$, this implies that
    $b_{0}\left\lVert Z\right\lVert_{H^{1}(B_{1}\setminus\mathcal{P})}^{2}=0.$
    thus implying that $b_{0}=0$.
  4. (4)
    Again, in view of (5.11) and (5.12) we have that
    $\frac{\widehat{Y}(x_{d-1},x_{d}\cos(\delta_{n})\pm x_{j}\sin(\delta_{n}))-\widehat{Y}(x_{d-1},x_{d})}{\delta_{n}}\to\mp\frac{3}{2}x_{j}V_{0}(x_{d-1},x_{d})$
    weakly in $H^{1}(\Omega_{i})$, for $i=1,2,3$, as $n\to\infty$. Passing to the limit in (5.13), reasoning as in step (2) and (3), we get
    $b_{j}\left\lVert x_{j}V_{0}\right\lVert_{H^{1}(B_{1}\setminus\mathcal{P})}^{2}=0.$
    which forces $b_{j}=0$, thus completing the proof of the claim.
Proof of Step 4. For any $n\in\mathbb{N}$, we define $u_{n}\in H^{1}(B_{1};\Sigma_{N_{n}})$ and $\widetilde{u}_{n,i}$, $i=1,2,3$, as in Step 1. Now, we compute the energy of $w_{n,i}$ and $\widetilde{u}_{n,i}$ on each $\Omega_{i}$ (setting $\Omega_{i}^{\prime}:=\partial\Omega_{i}\cap\partial B_{1}$, $\gamma:=3/2$ and extending $t=0$ for $r<1/2$), obtaining that
$\int_{\Omega_{i}}|\nabla w_{n,i}|^{2}\,\mathrm{d}x=\int_{0}^{1}r^{d-1}|\partial_{r}(r^{\gamma})|^{2}\,\mathrm{d}r\int_{\Omega_{i}^{\prime}}w_{n,i}^{2}\,\mathrm{d}S+\int_{0}^{1}r^{d+2\gamma-3}\,\mathrm{d}r\int_{\Omega_{i}^{\prime}}|\nabla_{\partial B_{1}}w_{n,i}|^{2}\,\mathrm{d}S$
and that
$\int_{\Omega_{i}}|\nabla\widetilde{u}_{n,i}|^{2}\,\mathrm{d}x$ $=\int_{\Omega_{i}\setminus B_{\frac{1}{2}}}|\nabla(tw_{n,i})|^{2}\,\mathrm{d}x$
$=\int_{0}^{1}r^{d-1}|\partial_{r}(tr^{\gamma})|^{2}\,\mathrm{d}r\int_{\Omega_{i}^{\prime}}w_{n,i}^{2}\,\mathrm{d}S+\int_{0}^{1}r^{d+2\gamma-3}t^{2}\,\mathrm{d}r\int_{\Omega_{i}^{\prime}}|\nabla_{\partial B_{1}}w_{n,i}|^{2}\,\mathrm{d}S.$
Plugging these computations into (5.5) and recalling (5.7), we get that
$\int_{0}^{1}r^{d+2\gamma-3}(1-t^{2})\,\mathrm{d}r\sum_{i=1}^{3}\int_{\Omega_{i}^{\prime}}|\nabla_{\partial B_{1}}w_{n,i}|^{2}\,\mathrm{d}S\\ <\int_{0}^{1}r^{d-1}\left(|\partial_{r}(tr^{\gamma})|^{2}-|\partial_{r}(r^{\gamma})|^{2}\right)\,\mathrm{d}r\sum_{i=1}^{3}\int_{\Omega_{i}^{\prime}}w_{n,i}^{2}\,\mathrm{d}S\\ \qquad-(d+1)\int_{0}^{1}r^{d}(1-\varepsilon_{n}-t)\,\mathrm{d}r\,\frac{\beta(c_{n},Y)}{\delta_{n}^{2}}+o(1),$
as $n\to\infty$. Since
$\int_{0}^{1}r^{d+2\gamma-3}(1-t^{2})\,\mathrm{d}r>0,\quad(d+1)\int_{0}^{1}r^{d}(1-\varepsilon_{n}-t)\,\mathrm{d}r\,\frac{\beta(c_{n},Y)}{\delta_{n}^{2}}\geq 0$
for $n$ sufficiently large and since, by compact trace embedding,
$w_{n,i}\to 0\quad\text{strongly in }L^{2}(\Omega_{i}^{\prime}),~{}\text{as }n\to\infty,$
we deduce that $w_{n,i}\to 0$ strongly in $H^{1}(\Omega_{i})$, for all $i=1,2,3$, thus reaching a contradiction and concluding the proof. ∎
6. Regularity of the free boundary around points of frequency $\nicefrac{{3}}{{2}}$
In this section we complete the proof of Theorem 1.1 by proving the regularity of the free interface $\mathcal{F}(u)$ around points of frequency $\nicefrac{{3}}{{2}}$. The main ingredients are the classification of the $3/2$-homogeneous blow-ups (Theorem 1.2) and the epiperimetric inequality Theorem 1.3.
6.1. Rate of convergence of the blow-up sequence
We will show that we can apply the epiperimetric inequality to the rescalings of $u$ of the form $u^{r,x_{0}}(x)=r^{-\nicefrac{{3}}{{2}}}u(x_{0}+rx)$. In order to do so, we need to show that the conditions of Theorem 1.3 are fulfilled uniformly, at all points $x_{0}$ and at all small scales $r$, by multiples of $u^{r,x_{0}}$. We start with the following
Lemma 6.1 (Uniform Hausdorff distance estimate).
Let $d\geq 2$ and $N\geq 1$ be fixed. For any $\varepsilon>0$ there exists $\delta>0$ (depending on $d$, $N$ and $\varepsilon$) such that, if $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ is a minimizer, $u\in\mathcal{M}(B_{1};N)$, that satisfies
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u,Y)\right\lVert_{L^{\infty}(B_{1})}\leq\delta,$
then
(6.1) $\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\Big{(}\{u_{i}>0\}\cap B_{\nicefrac{{1}}{{2}}},\{Y_{i}>0\}\cap B_{\nicefrac{{1}}{{2}}}\Big{)}\leq\varepsilon,$
and
(6.2) $\sum_{i=4}^{N}\mathrm{d}_{\mathcal{H}}\Big{(}\{u_{i}>0\}\cap B_{\nicefrac{{1}}{{2}}},\{x_{d-1}=x_{d}=0\}\cap B_{\nicefrac{{1}}{{2}}}\Big{)}\leq\varepsilon.$
Moreover, in $B_{1}\cap\left\{x\in\mathbb{R}^{d}\colon\sqrt{x_{d-1}^{2}+x_{d}^{2}}>\varepsilon\right\}$ the free boundary $\mathcal{F}(u)$ is the disjoint union of the three surfaces $\Gamma_{ij}=\partial\{u_{i}>0\}\cap\partial\{u_{j}>0\}$, for $i,j\in\{1,2,3\}$, $i\neq j$, which are $C^{1,\alpha}$ graphs over $\partial\{Y_{i}>0\}\cap\partial\{Y_{j}>0\}$.
Proof.
Suppose by contradiction that there are $\varepsilon>0$ and a sequence $u_{n}\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})\cap\mathcal{M}(B_{1};N)$ such that
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u_{n},Y)\right\lVert_{L^{\infty}(B_{1})}=\delta_{n}\to 0,$
and for which (6.1) or (6.2) fail. For any $\tau>0$, we denote
$\mathcal{O}_{\tau}$ :=$:=\{x\in\mathbb{R}^{d}\colon\operatorname{\mathrm{dist}}(x,\{Y=0\})<\tau\},$
$C_{\tau}$ :=$:=\left\{x\in\mathbb{R}^{d}\colon\sqrt{x_{d-1}^{2}+x_{d}^{2}}<\tau\right\}.$
Notice there exists $t=t_{\varepsilon}>0$ such that $\mathrm{d}_{\Sigma_{N}}(Y,0)>t$ on the set $B_{1}\setminus\mathcal{O}_{\varepsilon/6}.$ On the other hand, by the uniform convergence of $u_{n}$ to $Y$ we get that for $n$ large enough
$\{Y_{i}>0\}\cap B_{1}\setminus\mathcal{O}_{\varepsilon/6}\subset\{Y_{i}>t\}\cap B_{1}\subset\{u_{n,i}>0\}\cap B_{1}\quad\text{for}\quad i=1,2,3.$
In particular, this excludes the possibility that (6.1) fails for $u_{n}$ as $n\to+\infty$. In order to show that (6.2) cannot fail for large $n$, we notice that for every ball $B_{\varepsilon/3}(x)$ with center $x\in B_{\nicefrac{{1}}{{2}}}\cap\partial\{Y_{i}>0\}\cap\partial\{Y_{j}>0\}\setminus C_{\varepsilon/2}$ (for $i,j\in\{1,2,3\}$ with $i\neq j$), we have that $Y=Y_{i}\vec{\bm{e}}_{i}+Y_{j}\vec{\bm{e}}_{j}$ and
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u_{n},Y)\right\lVert_{L^{\infty}(B_{\varepsilon/3}(x))}\to 0\quad\text{as }n\to\infty.$
Thus, by the epsilon-regularity lemma [OV24, Lemma 9.8], we get that $u_{k}\equiv 0$ in $B_{\varepsilon/6}(x)$ for all $k\neq i,j$ and $\mathcal{F}(u)\cap B_{\varepsilon/6}(x)=\Gamma_{ij}\cap B_{\varepsilon/6}(x)$ is a $C^{1,\alpha}$ graph (with small $C^{1}$ norm) over the hyperplane $\partial\{Y_{i}>0\}\cap\partial\{Y_{j}>0\}$. In particular, this implies that $(2)$ cannot happen and it proves that second part of the claim. ∎
Lemma 6.2 (Uniform convergence of rescalings with variable center).
Let $u\in\mathcal{M}(D;N)$ be such that $\gamma(u,0)=\nicefrac{{3}}{{2}}$. Then, for all $\varepsilon>0$ there exists $\rho,r_{0}>0$ (depending on $d$, $N$ and $\varepsilon$) such that for all $z\in B_{\rho}\cap\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and for all $r<r_{0}$ there exists $Y\in\mathcal{M}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$ such that
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{z,r},Y)\right\lVert_{L^{\infty}(B_{\nicefrac{{1}}{{2}}})}\leq\varepsilon,$
where $u^{z,r}(x):=H(u,z,r)^{-\nicefrac{{1}}{{2}}}u(rx+z)$.
Proof.
Let us assume by contradiction that there exists sequences $r_{n}\to 0$ and $z_{n}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, with $|z_{n}|\to 0$, such that
(6.3) $\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{z_{n},r_{n}},Y)\right\lVert_{L^{\infty}(B_{\nicefrac{{1}}{{2}}})}>\varepsilon\quad\text{for all }n\in\mathbb{N}$
for all $Y\in\mathcal{M}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$, for some $\varepsilon>0$, and that $\gamma(u^{z_{n},r_{n}},0)=\nicefrac{{3}}{{2}}$ for all $n\in\mathbb{N}$. First of all, we observe that
(6.4) $\sum_{i=1}^{N}\int_{B_{1}}|\nabla u^{z_{n},r_{n}}_{i}|^{2}\,\mathrm{d}x=\mathcal{N}(u^{z_{n},r_{n}},0,1).$
Thus, up to a subsequence
(6.5) $u^{z_{n},r_{n}}\to u_{0}\quad\text{strongly in }H^{1}(B_{1};\Sigma_{N})~{}\text{and uniformly in }B_{\nicefrac{{1}}{{2}}},$
for some $u_{0}\in H^{1}(B_{1};\Sigma_{N})$. Moreover, since $u^{z_{n},r_{n}}\in\mathcal{M}(B_{1};N)$ for all $n$, also $u_{0}\in\mathcal{M}(B_{1};N)$. We now observe that, since $z_{n}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and by monotonicity, we have
$\frac{3}{2}=\mathcal{N}(u,z_{n},0^{+})\leq\mathcal{N}(u,z_{n},Rr_{n})=\mathcal{N}(u^{z_{n},r_{n}},0,R),$
for any $R>0$ and $n$ sufficiently large. Passing to the limit as $n\to\infty$, we get
$\frac{3}{2}\leq\mathcal{N}(u_{0},0,R)\quad\text{for all }R>0.$
By (6.5) and scaling properties, we have that
$\mathcal{N}(u_{0},0,R)=\lim_{n\to\infty}\mathcal{N}(u^{z_{n},r_{n}},0,R)=\lim_{n\to\infty}\mathcal{N}(u,z_{n},Rr_{n}),$
for all $R>0$. Therefore, by upper semicontinuity and monotonicity of the frequency function and Lemma 2.4, we have that for all $R,\tau>0$ sufficiently small, there holds
$\frac{3}{2}\leq\mathcal{N}(u_{0},0,R)\leq\mathcal{N}(u,0,\tau R).$
Sending $\tau\to 0$ implies that $u_{0}$ is $\nicefrac{{3}}{{2}}$-homogeneous, which, in view of Lemma 2.4 and (6.5) contradicts (6.3). ∎
We are now in position to prove the uniqueness of the blow-up limit and rate of convergence of the blow-up sequence.
Proposition 6.3.
Let $\alpha:=\varepsilon(d+1)$, with $\varepsilon>0$ being as in Theorem 1.3. Let $D\subseteq\mathbb{R}^{d}$ be open and let $u\in\mathcal{M}(D;N)$. Then, for all compact $K\subseteq D$ there exists $R>0$ such that for all $x_{0}\in K$ satisfying $\gamma(u,x_{0})=\nicefrac{{3}}{{2}}$, there exists $C_{\textup{rate}}>0$ depending on $d$ and $\mathcal{N}(u,x_{0},R)$ such that
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{x_{0},r_{2}},u^{x_{0},r_{1}})\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq\sum_{i=1}^{N}\left\lVert u_{i}^{x_{0},r_{2}}-u_{i}^{x_{0},r_{1}}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq C_{\textup{rate}}(r_{2}^{\alpha}-r_{1}^{\alpha})$
for all $0<r_{1}<r_{2}<R$ where $u^{x_{0},r}(x):=r^{-\nicefrac{{3}}{{2}}}u(rx+x_{0})$. In particular, there exists $Y^{x_{0}}\in\mathcal{M}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$ such that
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{x_{0},r},Y^{x_{0}})\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq\sum_{i=1}^{N}\left\lVert u_{i}^{x_{0},r}-Y^{x_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq C_{\textup{rate}}r^{\alpha},$
for all $r<\operatorname{\mathrm{dist}}(x_{0},\partial D)$.
Proof.
Thanks to Lemma 6.1 and Lemma 6.2, there is a radius $R$ such that we can apply the epiperimetric inequality to all the rescaling of the form $u^{x_{0},r}(x)=r^{-\nicefrac{{3}}{{2}}}u(x_{0}+rx)$ with $r\leq R$ and $x_{0}\in K$. The uniqueness of the blow-up and the rate of convergence now follow by a standard argument (see e.g. [OV24, Proposition 8.1]). ∎
As a corollary, we also obtain the following
Proposition 6.4 (Oscillation of the $\nicefrac{{3}}{{2}}$-blow-up limits).
Let $\bar{\alpha}:=2\alpha/(\alpha+3)$, where $\alpha$ is as in Proposition 6.3 and let $u\in\mathcal{M}(D;N)$. Then, for all compact $K\subseteq D$ there exists $r_{0}>0$ (depending on $d$ and $K$) and $C_{\textup{osc}}>0$ (depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$) such that
$\sum_{i=1}^{3}\left\lVert Y^{x_{0}}_{i}-Y^{z_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq C_{\textup{osc}}|x_{0}-z_{0}|^{\bar{\alpha}}$
for all $x_{0},z_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)\cap K$ such that $|x_{0}-z_{0}|<r_{0}$, where
$Y^{p}=\lim_{r\to 0}u^{p,r}\quad\text{and}\quad u^{p,r}(x):=r^{-\frac{3}{2}}u(rx+p).$
Proof.
We first estimate
(6.6) $\sum_{i=1}^{3}\left\lVert Y^{x_{0}}-Y^{z_{0}}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq 2\sum_{i=1}^{N}\Big{(}\left\lVert u^{x_{0},r}_{i}-Y^{x_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\\ +\left\lVert u^{z_{0},r}_{i}-Y^{z_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}+\left\lVert u^{x_{0},r}_{i}-u^{z_{0},r}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\Big{)}.$
The first two terms can be bounded in view of Proposition 6.3 as follows
(6.7) $\sum_{i=1}^{N}\left(\left\lVert u^{x_{0},r}_{i}-Y^{x_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}+\left\lVert u^{z_{0},r}_{i}-Y^{z_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\right)\leq C\,r^{\alpha},$
for $r$ sufficiently small depending on $d$ and $\mathrm{d}_{\mathcal{H}}(K,\partial D)$ and $C>0$ depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$. On the other hand, if $L_{K}>0$ denotes the Lipschitz constant of $u$ in $K$, we have that
(6.8) $\sum_{i=1}^{N}\left\lVert u^{x_{0},r}_{i}-u^{z_{0},r}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq\frac{L_{K}^{2}}{r^{3}}|x_{0}-z_{0}|^{2}.$
Therefore, we can choose $r$ in such a way that
$r^{\alpha}=\frac{|x_{0}-z_{0}|^{2}}{r^{3}},\quad\text{that is}\quad r=|x_{0}-z_{0}|^{\frac{2}{\alpha+3}}.$
The claim follows by combining (6.8) and (6.7) with (6.6). ∎
6.2. Regularity of $\mathcal{F}_{\nicefrac{{3}}{{2}}}$
In this section we conclude the proof of Theorem 1.1.
Lemma 6.5 (Frequency gap from above).
There exists $\delta_{d}>0$ such that, if $u\in\mathcal{M}_{\gamma}(\mathbb{R}^{d};N)$ is a $\gamma$-homogeneous minimizer with $\gamma>3/2$, then there holds $\gamma\geq 3/2+\delta_{d}$.
Proof.
Suppose that this is not the case and that there is a sequence of $\gamma_{n}$ homogeneous solutions $u_{n}$ with $\gamma_{n}\to\nicefrac{{3}}{{2}}$. Then, up to a subsequence, $u_{n}$ converges to a $\nicefrac{{3}}{{2}}$-homogeneous solution $Y$. By Lemma 6.1, for $n$ large enough we can apply the epiperimetric inequality from Theorem 1.3 to $u_{n}$. But this is a contradiction with the minimality of $u_{n}$. ∎
Lemma 6.6 (No holes lemma).
For any $\varepsilon,\rho\in(0,\nicefrac{{1}}{{2}})$ there exists $\delta>0$ (depending on $d$, $N$ and $\varepsilon$) such that, if $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})\cap\mathcal{M}(B_{1};N)$ satisfies
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u,Y)\right\lVert_{L^{\infty}(B_{1})}\leq\delta\quad\text{and}\quad\sup_{x\in B_{\frac{1}{2}}}\mathcal{N}\left(u,x,\frac{1}{2}\right)<\frac{3}{2}+\delta,$
then for all $x^{\prime\prime}\in B_{\rho}^{\prime\prime}:=\{y\in\mathbb{R}^{d-2}\colon|y|<\rho\}$ there exists $x=(x^{\prime\prime},x_{d-1},x_{d})$ such that
$\sqrt{x_{d-1}^{2}+x_{d}^{2}}<\varepsilon\qquad\text{and}\qquad\gamma(u,x)=\nicefrac{{3}}{{2}}.$
Proof.
We argue by contradiction and assume that there exist $\varepsilon,\rho\in(0,\nicefrac{{1}}{{2}})$, vanishing sequences $\{\delta_{n}\}\subseteq\mathbb{R}_{+}$ and $u_{n}\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u_{n},Y)\right\lVert_{L^{\infty}(B_{1})}\leq\delta_{n}\quad\text{and}\quad\sup_{x\in B_{\frac{1}{2}}}\mathcal{N}\left(u,x,\frac{1}{2}\right)<\frac{3}{2}+\delta_{n},$
for all $n\in\mathbb{N}$ and there exist $x^{\prime\prime}_{0}\in B_{\rho}^{\prime\prime}$ such that there holds $\gamma(u_{n},x)\neq\nicefrac{{3}}{{2}}$ for all points $x\in\mathcal{F}(u_{n})$ of the form $x=(x^{\prime\prime}_{0},x_{d-1},x_{d})\in B_{\rho}\cap C_{\varepsilon}$. In view of Lemma 6.5 and the fact that
$\gamma(u_{n},x)\leq\mathcal{N}\left(u_{n},x,\frac{1}{2}\right)<\frac{3}{2}+\delta_{n},$
for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$, we derive that $\gamma(u_{n},x)<\nicefrac{{3}}{{2}}$ for $n$ sufficiently large for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$. On the other hand, from Theorem 1.2 we obtain that $\gamma(u_{n},x)=1$ for all $x=(x^{\prime\prime}_{0},x_{d-1},x_{d})\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$. As a consequence, for some $n$ large enough, we can find $t>0$ such that in the set
$\mathcal{U}:=\Big{\{}(x^{\prime\prime},x_{d-1},x_{d})\ :\ |x^{\prime\prime}-x^{\prime\prime}_{0}|<t,\ \sqrt{x_{d-1}^{2}+x_{d}^{2}}<\frac{1}{2}\Big{\}},$
the whole free boundary $\mathcal{F}(u_{n})$ consists only of points of frequency $1$ and the number of connected components of $\mathcal{U}\setminus\mathcal{F}(u_{n})$ is finite. Let now $\omega$ be the connected component containing $B_{1/2}\cap\{u_{n,1}>0\}\setminus\mathcal{C}_{\varepsilon}$ and let $y\in B_{1/2}\cap\{u_{n,1}>0\}\setminus\mathcal{C}_{\varepsilon}$ be fixed. By Lemma 2.3 (applied to $\mathcal{M}=\mathcal{U}$ and $\mathcal{F}=\mathcal{F}(u_{n})$) every closed curve starting from $y$, contained in $\mathcal{U}$, and intersecting $\mathcal{F}(u_{n})$ a finite number of times (and in a transversal way) should cross $\mathcal{F}(u_{n})$ an even number of times. This is a contradiction since a closed curve circling around $\mathcal{C}_{\varepsilon}$ crosses $\mathcal{F}(u_{n})$ exactly three times, in view of Lemma 6.1. ∎
Proof of Theorem 1.1.
Let $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and let $u_{0}$ be the $\nicefrac{{3}}{{2}}$-homogeneous blow-up of $u$ at $x_{0}$, which is unique and non-zero by Proposition 6.3. Up to multiplying $u$ with a constant, we can suppose that $u_{0}=Y$. By Lemma 6.5 and Theorem 1.2 we know that in a neighborhood of $x_{0}$ there are only points of frequency $1$ or $\nicefrac{{3}}{{2}}$. By the no-holes lemma Lemma 6.6, we know that for every $x^{\prime\prime}\in\mathbb{R}^{d-2}$ with $|x^{\prime\prime}|$ small enough, there exist $(x_{d-1},x_{d})\in\mathbb{R}^{2}$ such that $x=(x^{\prime\prime},x_{d-1},x_{d})\in\nicefrac{{3}}{{2}}$. By Proposition 6.3, we know that the point $(x_{d-1},x_{d})$ is unique, so in a neighborhood of $x_{0}$, $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ is a graph of a function $\eta:\mathbb{R}^{d-2}\to\mathbb{R}^{2}$. The $C^{1,\alpha}$-regularity of $\eta$ follows from Proposition 6.3, Proposition 6.4 and Lemma 6.1 by a standard argument. Finally, the regularity of $\mathcal{F}(u)$ around $x_{0}$ is a consequence of the regularity of $\eta$, the uniqueness of the blow-up and the epsilon-regularity for $\mathcal{F}_{1}$ ([OV24, Lemma 9.8]); the argument is standard and for the details we refer for instance to [OV24, Section 9]. ∎
Acknowledgements
The authors are supported by the European Research Council (ERC), through the European Union’s Horizon 2020 project ERC VAREG - Variational approach to the regularity of the free boundaries (grant agreement No. 853404). The authors acknowledge the MIUR Excellence Department Project awarded to the Department of Mathematics, University of Pisa, CUP I57G22000700001. R.O. acknowledges support from the 2024 INdAM-GNAMPA project no. CUP_E53C23001670001. B.V. acknowledges support from the projects PRA_2022_14 GeoDom (PRA 2022 - Università di Pisa) and MUR-PRIN “NO3” (n.2022R537CS).

### v2

@@
 We consider energy-minimizing harmonic maps into trees and we prove the regularity of the singular part of the free interface near triple junction points.
-Precisely, by a dimension reduction and a topological argument, we prove that there are no admissible frequencies between $1$ and $\nicefrac{{3}}{{2}}$, and we classify the $\nicefrac{{3}}{{2}}$-homogeneous solutions.
-Then, by proving a new epiperimetric inequality, we show that around any point of frequency $\nicefrac{{3}}{{2}}$, the free interface is composed of three $C^{1,\alpha}$-smooth $(d-1)$-dimensional manifolds (of points of frequency $1$) with common $C^{1,\alpha}$-regular boundary (made of points of frequency $\nicefrac{{3}}{{2}}$) that meet along this boundary at 120 degree angles, which in particular implies that the free interface is a smooth deformation of the $Y$ cone.
+Precisely, by proving a new epiperimetric inequality, we show that around any point of frequency $\nicefrac{{3}}{{2}}$, the free interface is composed of three $C^{1,\alpha}$-smooth $(d-1)$-dimensional manifolds (composed of points of frequency $1$) with common $C^{1,\alpha}$-regular boundary (made of points of frequency $\nicefrac{{3}}{{2}}$) that meet along this boundary at 120 degree angles.
 Our results also apply to spectral optimal partition problems for the Dirichlet eigenvalues.
-Moreover, as a byproduct, we also prove that the $Y$-configuration is the only minimizer for the min-max optimal $3$-partition problem for the Dirichlet eigenvalues on the sphere, in any dimension.
 Key words and phrases:
@@
 We consider the model case in which the target space is of the form
-(1.1) $\Sigma_{N}:=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,,$
-By the works of Gromov-Schoen [GS92], Caffarelli-Lin [CL07, CL08], Conti-Terracini-Verzini [CTV02, CTV03, CTV05a] it is known that if
+(1.1) $\Sigma_{N}:=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,$
+endowed with the distance
+$\mathrm{d}_{\Sigma_{N}}(X,Y):=\begin{cases}|X_{i}-Y_{i}|,&\text{if }X_{j}=Y_{j}=0~{}\text{for all }j\neq i,\\ |X_{i}|+|Y_{j}|,&\text{if }X_{i},Y_{j}>0~{}\text{for some }i\neq j.\end{cases}$
+By the works of Gromov-Schoen [GS92], Caffarelli-Lin [CL07, CL08], Tavares-Terracini [TT12] and Soave-Terracini [ST15] it is known that if
 $\mathcal{F}(u)=\{x\in B_{1}\ :\ u(x)=0\},$
@@
 $E(u,B_{1}):=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x\,,$
-then $\mathcal{F}(u)$ can be decomposed into regular and singular parts according to the values of the Almgren’s frequency $\gamma(u,x)$ as follows:
-$\mathrm{Reg}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=1\big{\}}\qquad\text{and}\qquad\mathrm{Sing}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>1\big{\}}.$
-Gromov-Schoen and Caffarelli-Lin [GS92, CL08] studied the regular part $\mathrm{Reg}(u)$ in every dimension and showed that it is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92]) and is a $(d-2)$-rectifiable ([Alp20, Dee22]).
-Moreover, Conti, Terracini and Verzini showed in [CTV03] that the singular set in dimension $2$ is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point.
-In this paper we study the regularity of the singular set $\mathrm{Sing}(u)$ (and the structure of $\mathcal{F}(u)$ around singular points) in any dimension. Our main result is the following:
+then $\mathcal{F}(u)$ can be decomposed into regular and singular parts according to the values of the Almgren’s frequency $\gamma(u,x)$ (see Section 2.2) as follows:
+$\mathrm{Reg}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=1\big{\}}\qquad\text{and}\qquad\mathrm{Sing}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)\geq\nicefrac{{3}}{{2}}\big{\}}.$
+By [GS92, CL08], in every dimension $d\geq 2$, the regular part $\mathrm{Reg}(u)$ is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92]) and is $(d-2)$-rectifiable ([Alp20, Dee22]).
+In dimension $d=2$, Conti, Terracini and Verzini showed in [CTV03] that the singular set is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point (see also the recent paper [LMNS24] for the case of generic solutions).
+In this paper we study the regularity of the lowest stratum of $\mathrm{Sing}(u)$, that is
+(1.2) $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u):=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}},$
+and the structure of the entire free interface $\mathcal{F}(u)$ around singular points in $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ in any dimension $d\geq 2$. Our main result is the following:
 Theorem 1.1.
-Let $u:B_{1}\to\Sigma_{N}$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$ and $N\geq 1$. Then, the singular set can be decomposed as the disjoint union of
-$\mathrm{Sing}(u)=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
+Let $u:B_{1}\to\Sigma_{N}$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$ and $N\geq 1$.
+Then, the singular set of $\mathrm{Sing}{(u)}$ of the free interface $\mathcal{F}(u)$ can be decomposed as
+$\mathrm{Sing}{(u)}=\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)\cup\Big{\{}x\in\mathcal{F}{(u)}\ :\ \gamma(u,x)\geq\nicefrac{{3}}{{2}}+\delta_{d}\Big{\}},$
+where $\delta_{d}>0$ is dimensional constant.
+The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ is an open subset of $\mathrm{Sing}(u)$ and, locally, a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold.
+Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, the nodal set $\mathcal{F}(u)$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$.
+Results in the spirit of Theorem 1.1 exist in the context of area-minimizing currents, where the nodal set $\mathcal{F}(u)$ is replaced by the support of an area-minimizing current (mod $p=3$); in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93], in any dimension $d\geq 3$ (see also the lecture notes [Min24]); we also remark that the analysis for general co-dimension 1 area-minimizing currents mod p was completed only recently in [DHMS20], [DHM+21] and [DHM+22].
+To our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$.
+The main ingredient in the proof of Theorem 1.1 is the following epiperimetric inequality (Theorem 1.2) for the $\nicefrac{{3}}{{2}}$-Weiss energy
+$W_{\frac{3}{2}}(u):=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x-\frac{3}{2}\sum_{i=1}^{N}\int_{\partial B_{1}}u_{i}^{2}\,\mathrm{d}S$
+near homogeneous triple junctions $Y(x)=Y(x_{d-1},x_{d})$ with $2$-dimensional profile defined as
+(1.3) $Y_{i}(r,\theta)=\begin{cases}r^{\frac{3}{2}}\left|\cos\left(\frac{3}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{3}(i-1)\leq\theta\leq\pi-\frac{2\pi}{3}i,\\ 0,&\text{elsewhere},\end{cases}$
+for $i=1,2,3$.
+Theorem 1.2 (Epiperimetric inequality).
+There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on the dimension $d$ such that the following holds.
+For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that
+$\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{Y_{i}>0\}\right)+\sum_{i=4}^{N}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{x_{d-1}=x_{d}=0\}\right)\leq\tau$
+and
+$\left\lVert\mathrm{d}_{\Sigma_{N}}(c,Y)\right\lVert_{H^{1}(B_{1})}\leq\delta,$
+there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u=c$ on $\partial B_{1}$ and
+$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
+The proof of Theorem 1.2 is obtained via a contradiction argument in the spirit of Weiss [Wei99] (see also [FS16, GPS16]).
+The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, new in this field.
+1.1. Two consequences of Theorem 1.1
+We now discuss two corollaries of Theorem 1.1 for harmonic maps with values in metric graphs and for spectral optimal partitions.
+First of all we notice that an immediate corollary of Theorem 1.1 is that the same result holds for harmonic maps with values in a locally finite tree $\Sigma$.
+Indeed, by [GS92, Dee22], if $u:B_{1}\to\Sigma$ is an energy-minimizing map defined in $B_{1}\subset\mathbb{R}^{d}$ with values in a locally finite metric tree, and if $p$ is a vertex of $\Sigma$, then the set $\mathcal{F}{(u)}:=u^{-1}(\{p\})$ can be decomposed into a regular and a singular parts, $\mathrm{Reg}{(u)}$ and $\mathrm{Sing}{(u)}$, where $\mathrm{Reg}{(u)}$ is a a smooth $(d-1)$-manifold that consists of the points of order (frequency) ${\rm Ord}^{u}=1$, while $\mathrm{Sing}{(u)}$ is $(d-2)$-rectifiable and consists of of points of order (frequency) ${\rm Ord}^{u}>1$; furthermore, for every $x\in\mathcal{F}{(u)}$, there is a neighborhood $\Omega\subset B_{1}$ such that $u:\Omega\to\Sigma$ has values in a subtree $\Sigma_{N}$ of the form (1.1).
+Thus, we have the following.
+Theorem 1.3.
+Let $\Sigma$ be a metric tree which is locally finite and let $u:B_{1}\to\Sigma$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$.
+Then, for every vertex $p$ of $\Sigma$ the singular set of $\mathcal{F}{(u)}:=u^{-1}(\{p\})$ can be decomposed as
+$\mathrm{Sing}{(u)}=\Big{\{}x\in\mathcal{F}(u)\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ {\rm Ord}^{u}(x)\geq\nicefrac{{3}}{{2}}+\delta_{d}\Big{\}},$
+where $\delta_{d}>0$ is a dimensional constant.
+The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}:=\big{\{}x\in\mathcal{F}{(u)}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold.
+Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, the set $\mathcal{F}{(u)}$ consists of three $(d-1)$-manifolds with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$.
+Another consequence concerns the regularity of the following optimal partition problem studied in [CL07], [CTV03, CTV05a] and [OV24].
+Given a bounded open set $D\subset\mathbb{R}^{d}$, $d\geq 2$, and a natural number $N\geq 2$ we consider the optimal partition problem
+(1.4) $\min\Big{\{}\sum_{j=1}^{N}\lambda_{1}(\Omega_{j})\ :\ \Omega_{j}\,-\,\text{open},\ \Omega_{j}\subset D,\ \Omega_{i}\cap\Omega_{j}=\emptyset\ \text{for}\ i\neq j\Big{\}},$
+where $\lambda_{1}(\Omega_{j})$ is the first eigenvalue on $\Omega_{j}$ with Dirichlet boundary conditions on $\partial\Omega_{j}$.
+If the $N$-uple of disjoint sets $(\Omega_{1},\dots,\Omega_{N})$ is a minimizer of (1.4) and if $u=(u_{1},\dots,u_{N})$ is the vector of the normalized first eigenfunctions (each $u_{j}$ is extended by zero outside $\Omega_{j}$), then the free interface
+$\mathcal{F}{(u)}:=D\cap\Big{(}{\bigcup_{j=1}^{N}}\partial\Omega_{j}\Big{)}$
+can be decomposed as a regular part and a singular part as follows (see [CL07, ST15, Alp20]): the regular part $\mathrm{Reg}{(u)}$ consists of all points $x\in\mathcal{F}{(u)}$ of frequency $\gamma(u,x)=1$ and is a smooth manifold of dimension $d-1$; the singular part $\mathrm{Sing}{(u)}$ is a closed $(d-2)$-rectifiable set composed of all the points of frequency $\gamma(u,x)\geq\nicefrac{{3}}{{2}}$.
+We have the following theorem.
+Theorem 1.4.
+Let $(\Omega_{1},\dots,\Omega_{N})$ be a minimizer of (1.4) and $u=(u_{1},\dots,u_{N})$ be the corresponding vector of the normalized first eigenfunctions.
+Then, the singular set of $\mathrm{Sing}{(u)}$ of the free interface $\mathcal{F}(u)$ can be decomposed as
+$\mathrm{Sing}{(u)}=\Big{\{}x\in\mathcal{F}{(u)}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}{(u)}\ :\ \gamma(u,x)\geq\nicefrac{{3}}{{2}}+\delta_{d}\Big{\}},$
+where $\delta_{d}>0$ is dimensional constant.
 The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}:=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold.
-Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, the nodal set $\mathcal{F}(u)$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$.
-The main ingredients for the proof of Theorem 1.1 are the classification of the singular homogeneous solutions of lowest frequency, which we achieve in Theorem 1.2 via a topological lemma combined with a dimension reduction argument (see Section 2) and a new epiperimetric inequality (Theorem 1.3) which we obtain via a contradiction argument in the spirit of Weiss [Wei99] (see also [GPS16, FS16]).
-Furthermore, Theorem 1.1 solves the problem of finding the optimal frequency gap in every dimension $d\geq 2$.
-Using this result, in Theorem 1.5 we also solve, in any dimension, an open problem by Helffer, Hoffmann-Ostenhof and Terracini in [HHT09], which is the limit case of the Bishop-Friedland-Hayman’s conjecture [Bis92, FH76].
-Finally, in Theorem 1.6 and Theorem 1.7, we discuss two corollaries of Theorem 1.1 in the context of harmonic maps and optimal partition problems.
-Results in the spirit of Theorem 1.1 exist in the context of area-minimizing currents, where the nodal set $\mathcal{F}(u)$ is replaced by the support of an area-minimizing current (mod $p=3$); in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93], in any dimension $d\geq 3$ (see also the lecture notes [Min24]); we also remark that the analysis for general co-dimension 1 area-minimizing currents mod p was completed only recently in [DHMS20], [DHM+21] and [DHM+22].
-To our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$; we also notice that in the framework of Theorem 1.1 the classification of the tangent maps was (until now) an open problem even in dimension $d=3$ and cannot be obtained via elementary geometric arguments.
-The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, also new in this field.
-In the rest of the introduction we give a detailed account on the known results (Section 1.1); in Section 1.2 we state our optimal gap result (Theorem 1.2) and the epiperimetric inequality (Theorem 1.3); finally, Section 1.3 is dedicated to the Helffer-Hoffmann-Ostenhof-Terracini’s problem and the Bishop-Friedland-Hayman’s conjecture.
-1.1. Notations and known results
-Let us fix the dimension $d\geq 2$.
-We work with a general open set $D\subseteq\mathbb{R}^{d}$ and with functions $u=(u_{1},\dots,u_{N})\colon D\to\mathbb{R}^{N}$ with $N\geq 2$ components whose support in $D$ is disjoint, i.e.
-$u_{i}u_{j}\equiv 0$ in $D$.
-More precisely, we consider the singular (non-positively curved) target space
-$\Sigma_{N}:=\Bigg{\{}X\in\mathbb{R}^{N}\colon\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N}X_{i}^{2}X_{j}^{2}=0\Bigg{\}}$ =$=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,,$
-on which it is naturally defined the distance
-$\mathrm{d}_{\Sigma_{N}}(X,Y):=\begin{cases}|X_{i}-Y_{i}|,&\text{if }X_{j}=Y_{j}=0~{}\text{for all }j\neq i,\\ |X_{i}|+|Y_{j}|,&\text{if }X_{i},Y_{j}>0~{}\text{for some }i\neq j.\end{cases}$
-We then consider the Sobolev-type space
-$H^{1}(D;\Sigma_{N})=\left\{u\in(H^{1}(D))^{N}\colon u_{i}u_{j}=0~{}\text{a.e. in }D~{}\text{for all }i\neq j\right\}.$
-The space $H^{1}_{\textup{loc}}(D;\Sigma_{N})$ is analogously defined.
+Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, the set $\mathcal{F}{(u)}$ consists of three $(d-1)$-manifolds with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$.
+The rest of the paper is organized as follows.
+In Section 2 we fix the notation, the functional setting and we recall some known facts about segregated solutions.
+In Section 3, we classify the $\nicefrac{{3}}{{2}}$-homogeneous solutions and we find, as a corollary, the shape of minimizers of a certain optimal partition problem on the sphere (see Corollary 3.6).
+Section 5 is devoted to the proof of our main tool, the epiperimetric inequality Theorem 1.2 and, finally, in Section 6 we conclude the proof of our main regularity result Theorem 1.1.
+2. Preliminaries
+In this section we introduce the functional setting and we define the main tools in our framework, such as monotonicity formulas.
+For any $N\geq 1$, let $\Sigma_{N}$ be as in (1.1). For any open set $D\subseteq\mathbb{R}^{d}$, we define the Sobolev space
+$H^{1}(D;\Sigma_{N})=\left\{u\in(H^{1}(D))^{N}\colon u_{i}u_{j}=0~{}\text{a.e. in }D~{}\text{for all }i\neq j\right\},$
+with the space $H^{1}_{\textup{loc}}(D;\Sigma_{N})$ being analogously defined; we refer to [KS93] for the theory of Sobolev spaces in this setting.
 Now, to each open $D\subseteq\mathbb{R}^{d}$ and each function $u\in H^{1}(D;\Sigma_{N})$ we associate the corresponding Dirichlet energy, defined as
@@
 for all open $\Omega\Subset D$ and all $v\in H^{1}(\Omega;\Sigma_{N})$ such that $u-v\in H^{1}_{0}(\Omega;\Sigma_{N})$.
-On the other hand, some of our results hold true in a larger class of “critical points”, namely, the so called class $\mathcal{S}$.
-In order to define it, we first introduce some notation which is also useful in the rest of the paper.
-Given $u\in\mathcal{A}(D;N)$ and a function
-$a\colon\{1,\dots,N\}\to\{-1,1\}$
-we define the real-valued function $a\cdot u\in H^{1}_{\textup{loc}}(D)$ as
-$(a\cdot u)(x):=\sum_{i=1}^{N}a(i)u_{i}(x)\quad\text{for }x\in D.$
-Moreover, for any $i\in\{1,\dots,N\}$, we denote by $\sigma_{i}:=\{1,\dots,N\}\to\{-1,1\}$ the function
-$\sigma_{i}(k):=\begin{cases}1,&\text{if }k=i,\\ -1,&\text{if }k\neq 1,\end{cases}$
-so that
-$\sigma_{i}\cdot u=u_{i}-\sum_{j\neq i}u_{j}\quad\text{in $D$, for all }i=1,\dots,N.$
-Now, we say that a function $u\in\mathcal{A}(D;N)$ belongs to the class $\mathcal{S}(D;N)$ if
-$-\Delta u_{i}&\leq 0,&&\text{in }D,\\ -\Delta(\sigma_{i}\cdot u)&\geq 0,&&\text{in }D,\end{aligned}\right.$
-in the sense of distributions, for all $i=1,\dots,N$. In particular, there holds $\mathcal{M}(D;N)\subseteq\mathcal{S}(D;N)$ (see e.g. [CTV05c, Theorem 5.1]).
-The aim of present paper is to investigate regularity questions for solutions $u$ in the classes $\mathcal{S}(D;N)$ and $\mathcal{M}(D;N)$ focusing, in particular, on the geometrical structure and regularity of the free boundary, defined as
-$\mathcal{F}(u):=\bigcup_{i=1}^{N}\partial\{x\in D\colon u_{i}(x)>0\},$
-in the interior of $D$. We point out that $u\in C^{0,1}_{\textup{loc}}(D;\mathbb{R}^{N})$, see e.g. [CTV05c], which implies that the sets
+We also introduce the class $\mathcal{S}(D;N)$ of functions $u\in\mathcal{A}(D;N)$ satisfying
+$-\Delta u_{i}&\leq 0,&&\text{in }D,\\ -\Delta\Big{(}u_{i}-\sum_{j\neq i}u_{j}\Big{)}&\geq 0,&&\text{in }D,\end{aligned}\right.$
+in the sense of distributions, for all $i=1,\dots,N$. In particular, there holds $\mathcal{M}(D;N)\subseteq\mathcal{S}(D;N)$ (see e.g. [CTV05b, Theorem 5.1]).
+We now recall the main basic properties of critical points $u\in\mathcal{S}(D;N)$ and energy-minimizers $u\in\mathcal{M}(D;N)$ and of the free boundary $\mathcal{F}(u)$.
+Most of the properties we recall were proved in a series of papers, namely [GS92, CL07] for the class $\mathcal{M}(D;N)$ and [CTV05b, TT12, ST15] for $\mathcal{S}(D;N)$; for what concerns the regularity of different classes of segregated solutions we refer to [NRS24, ST24b, ST24a].
+2.1. Lipschitz continuity
+If $u\in\mathcal{S}(D;N)$, then it is locally Lipschitz continuous, i.e. $u\in C^{0,1}_{\textup{loc}}(D;\mathbb{R}^{N})$, which implies that the sets
 $\Omega^{u}_{i}:=\{x\in D\colon u_{i}(x)>0\}$
-are open and well-defined.
-Before stating our main results in Section 1.2, from Section 1.1.1 to Section 1.1.6 we review some known facts concerning critical points $u\in\mathcal{S}(D;N)$.
-Most of the properties we recall were proved in a series of papers, namely [GS92, Sun03, CL07, Alp20] for the class $\mathcal{M}(D;N)$ and [CTV05c, TT12] for $\mathcal{S}(D;N)$.
-1.1.1. Almgren’s frequency function
-First of all, for any non-trivial $u\in\mathcal{S}(D;N)$ one can define the Almgren frequency function for any $x_{0}\in D$ and any $r<\operatorname{\mathrm{dist}}(x_{0},\partial D)$ as
-$\mathcal{N}(u,r,x_{0}):=\frac{E(u,r,x_{0})}{H(u,r,x_{0})},$
+are open and well-defined and that $u_{i}$ are harmonic on $\Omega_{i}^{u}$.
+2.2. Almgren and Weiss monotonicity formulas
+For any non-trivial $u\in\mathcal{S}(D;N)$ one can define the Almgren frequency function for any $x_{0}\in D$ and any $r<\operatorname{\mathrm{dist}}(x_{0},\partial D)$ as
+$\mathcal{N}(u,x_{0},r):=\frac{E(u,x_{0},r)}{H(u,x_{0},r)},$
 where
-$E(u,r,x_{0}):=\frac{1}{r^{d-2}}E(u,B_{r}(x_{0}))=\frac{1}{r^{d-2}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}|\nabla u_{i}|^{2}\,\mathrm{d}x$
+$E(u,x_{0},r):=\frac{1}{r^{d-2}}E(u,B_{r}(x_{0}))=\frac{1}{r^{d-2}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}|\nabla u_{i}|^{2}\,\mathrm{d}x$
 is the scaled energy and
-$H(u,r,x_{0}):=\frac{1}{r^{d-1}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}u_{i}^{2}\,\mathrm{d}S$
-is the scaled height function. The crucial fact is that the $r\mapsto\mathcal{N}(u,r,x_{0})$ is monotone non-decreasing and so the frequency of $u$ at the point $x_{0}$
-$\gamma(u,x_{0}):=\lim_{r\to 0^{+}}\mathcal{N}(u,r,x_{0})$
-is well-defined. As a consequence, one can derive unique continuation properties implying that the nodal set
-$\mathcal{Z}(u)=\{x\in D\colon u_{i}(x)=0~{}\text{for all }i=1,\dots,N\}$
-has empty interior and coincides with the free interface, i.e. $\mathcal{Z}(u)=\mathcal{F}(u)$. Moreover, there holds
-$\text{either}\quad\gamma(u,x_{0})=1\quad\text{or}\quad\gamma(u,x_{0})\geq 1+\delta_{d},$
-for all $x_{0}\in\mathcal{F}(u)$, with $\delta_{d}>0$ being a dimensional constant. Once defined, for any $\gamma\geq 1$,
-$\mathcal{F}_{\gamma}(u):=\{x\in\mathcal{F}(u)\colon\gamma(u,x_{0})=\gamma\},$
-we split
-$\mathcal{F}(u)=\mathrm{Reg}(u)\cup\mathrm{Sing}(u),$
-where
-$\mathrm{Reg}(u):=\mathcal{F}_{1}(u)\quad\text{and}\quad\mathrm{Sing}(u):=\bigcup_{\gamma\geq 1+\delta_{d}}\mathcal{F}_{\gamma}(u),$
-with $\mathrm{Reg}(u)$ relatively open and $\mathrm{Sing}(u)$ relatively closed.
-1.1.2. Blow-up analysis
+$H(u,x_{0},r):=\frac{1}{r^{d-1}}\sum_{i=1}^{N}\int_{B_{r}(x_{0})}u_{i}^{2}\,\mathrm{d}S$
+is the scaled height function. It is known that for $u\in\mathcal{S}(D;N)$ the function $r\mapsto\mathcal{N}(u,x_{0},r)$ is monotone non-decreasing and so the frequency of $u$ at $x_{0}$
+$\gamma(u,x_{0}):=\lim_{r\to 0^{+}}\mathcal{N}(u,x_{0},r)$
+is well-defined at any point $x_{0}\in\mathcal{F}(u)$. As a consequence, one can derive unique continuation properties implying that $\mathcal{F}(u)$ has empty interior and that
+$\mathcal{F}(u)=\bigcup_{i=1}^{N}\partial\{x\in D\colon u_{i}(x)>0\}.$
+Hence, we can also split the points of the free boundary in terms of their frequency; namely, for any $\gamma>0$ we let
+$\mathcal{F}_{\gamma}(u):=\{x\in\mathcal{F}(u)\colon\gamma(u,x)=\gamma\}.$
+Finally, we point out that also the Weiss energy
+$W_{\gamma}(u,x_{0},r):=\frac{H(u,x_{0},r)}{r^{2\gamma}}\Big{(}\mathcal{N}(u,x_{0},r)-\gamma\Big{)}$
+is monotone non-decreasing with respect to $r>0$, for any $\gamma>0$.
+2.3. Blow-up analysis
 Let $u\in\mathcal{S}(D;N)$, respectively $\mathcal{M}(D;N)$, and let $x_{0}\in\mathcal{F}_{\gamma}(u)$, for some $\gamma\geq 1$. We define the Almgren rescalings as
@@
 […]
-1.1.3. Points of frequency 1
-Let $u\in\mathcal{S}(N,D)$ and let $x_{0}\in\mathcal{F}_{1}(u)$. Then, the limit $U\in\mathcal{S}_{1}(\mathbb{R}^{d};N)$ is unique (i.e. independent from the chosen subsequence) and of the form
-$U_{1}=x_{d}^{-},\quad U_{2}=x_{d}^{+},\quad U_{i}\equiv 0~{}\text{for }i=3,\dots,N,$
-up to rotations and dilations and there holds
-$\frac{u(rx+x_{0})}{r}\to a\,U,~{}\text{uniformly in $B_{R}$ and in $H^{1}(B_{R};\mathbb{R}^{N})$, as }r\to 0,$
-for some $a\neq 0$ and all $R>0$.
-Moreover, around $x_{0}$, the free interface $\mathcal{F}(u)$ is a smooth manifold separating two of the sets $\Omega_{i}^{u}$.
-Precisely, there is a neighborhood $B_{r}(x_{0})$ and two different indices $i,j\in\{1,\dots,N\}$ such that
-$\Omega_{k}^{u}\cap B_{r}(x_{0})=\emptyset\quad\text{for every}\quad k\notin\{i,j\},$
-while in $B_{r}(x_{0})$ the function $u_{i}-u_{j}$ is harmonic and
-$|\nabla(u_{i}-u_{j})(x_{0})|\neq 0.$
-In other words, by choosing $r$ small enough, the interface
-$\{u_{i}-u_{j}=0\}\cap B_{r}(x_{0}),$
-is smooth $(d-1)$-manifold and we have
-$\mathcal{F}(u)\cap B_{r}(x_{0})=\partial\Omega_{i}^{u}\cap B_{r}(x_{0})=\partial\Omega_{j}^{u}\cap B_{r}(x_{0})=\{u_{i}-u_{j}=0\}\cap B_{r}(x_{0}).$
-The same statements are true replacing $\mathcal{S}(D;N)$ with $\mathcal{M}(D;N)$ and $\mathcal{S}_{1}(\mathbb{R}^{d};N)$ with $\mathcal{M}_{1}(\mathbb{R}^{d};N)$ (which, actually, coincides with $\mathcal{S}_{1}(\mathbb{R}^{d};N)$).
-1.1.4. Classification of blow-ups in dimension 2
-In dimension $d=2$, the admissible frequencies and the blow-up limits are completely classified. Namely, it is known that if $x_{0}\in\mathcal{F}(u)$, then
-$\gamma(u,x_{0})=\frac{m}{2}\quad\text{for some }m\in\mathbb{N},\ m\geq 2\,.$
-Moreover, the sets $\mathcal{S}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ and $\mathcal{M}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ coincide and they contain a unique (up to rotations and dilations) segregated, $\nicefrac{{m}}{{2}}$-homogeneous function.
-More precisely, if we assume that all the $N$ components are non-trivial, then $N=m$ and any blow-up limit $u\in\mathcal{S}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)=\mathcal{M}_{\nicefrac{{m}}{{2}}}(\mathbb{R}^{d};N)$ is of the form
-$u_{i}(r,\theta)=\begin{cases}r^{\frac{m}{2}}\left|\cos\left(\frac{m}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{m}(i-1)\leq\theta\leq\pi-\frac{2\pi}{m}i,\\ 0,&\text{elsewhere},\end{cases}$
-up to rotations and dilations, for $i=1,\dots,N$.
-1.1.5. Clean-up
-In [Sun03], the author works in the framework of harmonic maps with $\mathbb{R}$-trees as target space (such as $\Sigma_{N}$).
-More precisely, in [Sun03, Theorem 1.2] it is proved what, in the language of optimal partition problems, is known as a clean-up result.
-Once translated in our framework, it says that for any $\gamma\geq 1$ there exists $\varepsilon,\rho>0$ depending on $d$ and $\gamma$ such that if $u\in\mathcal{M}(D;N)$ satisfies
-$\mathcal{N}(u,r,x_{0})\leq\gamma\quad\text{and}\quad\left\lVert u_{i}\right\lVert_{L^{\infty}(B_{r}(x_{0}))}\leq\varepsilon\sqrt{H(u,r,x_{0})},$
-for some $x_{0}\in D$, $r>0$ (such that $\overline{B_{r}(x_{0})}\subseteq D$) and $i\in I$, with $I\subseteq\{1,\dots,N\}$, then
-$u_{i}\equiv 0\quad\text{in }B_{r\rho}(x_{0})\quad\text{for all }i\in I.$
-In broad terms, if some components of a minimizer are small in a ball, then they disappear in a smaller ball.
-1.1.6. The singular set
-By a dimension reduction argument, for any $u\in\mathcal{M}(D;N)$ we have
-$\dim_{\mathcal{H}}\mathrm{Sing}(u)\leq d-2.$
-Furthermore, thanks to the techniques introduced in [NV17, NV18], in [Alp20, Dee22] it was shown that $\mathrm{Sing}(u)$ is $(d-2)$-rectifiable and that satisfies the following estimate
-$|B_{r}(\mathrm{Sing}(u)\cap K)|\leq C(K,u)r^{2},$
-for all compact $K\subseteq D$ and all $r>0$ sufficiently small.
-1.2. New results on the structure of the singular set
-Our main result (Theorem 1.1) provides a fine description of the singular set and of the whole nodal set around points in the lowest stratum.
-The first question we address is the detection of the optimal frequency gap $\delta_{d}$, defined in Section 1.1.1; until now, the frequency gap constant was explicitly known only in dimension $2$, i.e.
-$\delta_{2}=\nicefrac{{1}}{{2}}$ (see Section 1.1.4).
-Hence, our first main result extends this property to any dimension $d\geq 2$.
-We point out that this is true not only for local minimizers, but for all functions in the class $\mathcal{S}(D;N)$.
-Theorem 1.2 (Optimal gap).
-Let $u\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ be a non-trivial, $\gamma$-homogeneous critical point, with $\gamma\geq 1$. Then, either
-$\gamma=1\quad\text{and}\quad N=2$
-or
-$\gamma\geq\frac{3}{2}\quad\text{and}\quad N\geq 3.$
-Moreover, if $\gamma=\nicefrac{{3}}{{2}}$, then $N=3$ and, up to rotations and dilations, $u$ takes the form of a cylindrical $Y$-configuration, that is
-$u(x_{1},\dots,x_{d})=Y(x_{d-1},x_{d}),$
-where $Y=(Y_{1},Y_{2},Y_{3})\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{2};3)$ is defined as
-$Y_{i}(r,\theta)=\begin{cases}r^{\frac{3}{2}}\left|\cos\left(\frac{3}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{3}(i-1)\leq\theta\leq\pi-\frac{2\pi}{3}i,\\ 0,&\text{elsewhere},\end{cases}$
-for $i=1,2,3$.
-The above theorem proves that for any $u\in\mathcal{S}(D,N)$ (and, in particular, for any minimizer $u\in\mathcal{M}(D,N)$) the singular set can be decomposed as follows
-$\mathrm{Sing}(u)=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
-Moreover, every blow-up limit of $u$ at a point $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ is of the form $Y$. This, of course, does not imply that the set
-$\mathcal{F}_{\nicefrac{{3}}{{2}}}(u):=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}$
-is smooth around $x_{0}$ and not even that the blow-up at $x_{0}$ is unique (as it might depend on the sequence of rescalings $u_{r_{n},x_{0}}$).
-We next focus on the regularity and the geometrical structure of set $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$.
-Our approach in tackling this question is through an epiperimetric inequality for $\nicefrac{{3}}{{2}}$-Weiss energy, defined as
-$W_{\frac{3}{2}}(u):=E(u,1,0)-\frac{3}{2}H(u,1,0)=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x-\frac{3}{2}\sum_{i=1}^{N}\int_{\partial B_{1}}u_{i}^{2}\,\mathrm{d}S,$
-for $u\in H^{1}(B_{1};\Sigma_{N})$. More precisely, we have the following.
-Theorem 1.3 (Epiperimetric inequality).
-There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on the dimension $d$ such that the following holds.
-For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that
-$\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{Y_{i}>0\}\right)+\sum_{i=4}^{N}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{x_{d-1}=x_{d}=0\}\right)\leq\tau$
-and
-$\left\lVert\mathrm{d}_{\Sigma_{N}}(c,Y)\right\lVert_{H^{1}(B_{1})}\leq\delta,$
-there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u=c$ on $\partial B_{1}$ and
-$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
-We prove Theorem 1.2 in Section 2. Then, in Section 4 and Section 5 we prove Theorem 1.3. Finally, using this epiperimetric inequality, in Section 6 we prove Theorem 1.1.
-1.3. The Bishop-Friedland-Hayman’s conjecture for the min-max
-In this section we use Theorem 1.2 to show that the optimal min-max $3$-partition of the sphere is given by the triple junction configuration in any dimension.
-For any $p\in(0,+\infty]$ we consider the following family of optimal partition problems studied by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10]
-($\text{OP}_{p}$) $\min\left\{\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t.
-}\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\},$
-where $\lambda_{1}(\omega)$ denotes the first Dirichlet eigenvalue of the spherical Laplacian in $\omega\subseteq\partial B_{1}$, and where in the limit case $p=+\infty$, the value
-$\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}$
-is intended as
-$\max_{i\in\{1,\dots,N\}}\lambda_{1}(\omega_{i}).$
-In the particular case of the $3$-partition, i.e. $N=3$, we have the following conjecture.
-Problem 1.4 (Bishop-Friedland-Hayman-Helffer-Hoffmann-Ostenhof-Terracini).
-For $N=3$, prove that the $Y$-configuration is the unique minimizer of ($\text{OP}_{p}$), for any $p\in(0,+\infty]$.
-This conjecture was originally posed in [Bis92] in the case $p=1$, but can be generalized for any $p\in(0,+\infty]$ (see [FH76] and [HHT10]).
-Of particular interest is the case $p=1/2$, which is related with the best exponent in the multi-phase Alt-Caffarelli-Friedman monotonicity formula; in general, there is a vast literature on these optimal partition problems also in relation to models in biology and dynamics of populations, we refer for instance to [CTV03, CTV05b, BBH98, CL07, HHT09, HHT10].
-1.4 was proved for $p=+\infty$ in dimension $d=3$ by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10]; in Section 3 we prove the following theorem, which solves 1.4 (for $p=+\infty$) in any dimension $d\geq 3$.
-Theorem 1.5.
-The $Y$-configuration is the unique minimizer of
-$\min\left\{\max_{i\in\{1,2,3\}}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\}.$
-1.4. Two consequences of Theorem 1.1
-In this section we discuss two corollaries of Theorem 1.1 for harmonic maps with values in metric graphs and for spectral optimal partitions.
-First of all we notice that an immediate corollary of Theorem 1.1 (and the clean-up theorem from [Sun03]), is that the same result holds for harmonic maps with values in a locally finite tree $\Sigma$.
-Indeed, by [GS92, Sun03, Dee22], if $u:B_{1}\to\Sigma$ is an energy-minimizing map defined in $B_{1}\subset\mathbb{R}^{d}$ with values in a locally finite metric tree, and if $p$ is a vertex of $\Sigma$, then the set $\mathcal{F}:=u^{-1}(\{p\})$ can be decomposed into a regular and a singular parts, $Reg$ and $Sing$, where $Reg$ is a a smooth $(d-1)$-manifold that consists of the points of order (frequency) ${\rm Ord}^{u}$, while $Sing$ has Hausdorff dimension $(d-2)$ and consists of of points of order (frequency) ${\rm Ord}^{u}>1$; furthermore, for every $x\in\mathcal{F}$, there is a neighborhood $\Omega\subset B_{1}$ such that $u:\Omega\to\Sigma$ has values in a subtree $\Sigma_{N}$ of the form (1.1).
-Thus, we have the following
-Theorem 1.6.
-Let $\Sigma$ be a metric tree which is locally finite and let $u:B_{1}\to\Sigma$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$.
-Then, for every vertex $p$ of $\Sigma$ the singular set of $\mathcal{F}:=u^{-1}(\{p\})$ can be decomposed as the disjoint union of
-$\mathrm{Sing}=\Big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)>\nicefrac{{3}}{{2}}\Big{\}}.$
-The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}:=\big{\{}x\in\mathcal{F}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold.
-Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$, the set $\mathcal{F}$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}$.
-Another consequence concerns the regularity of the following optimal partition problem studied in [CL07], [CTV03] and [OV24].
-Given a bounded open set $D\subset\mathbb{R}^{d}$, $d\geq 2$, and a natural number $N\geq 2$ we consider the optimal partition problem
-(1.2) $\min\Big{\{}\sum_{j=1}^{N}\lambda_{1}(\Omega_{j})\ :\ \Omega_{j}\,-\,\text{open},\ \Omega_{j}\subset D,\ \Omega_{i}\cap\Omega_{j}=\emptyset\ \text{for}\ i\neq j\Big{\}},$
-where $\lambda_{1}(\Omega_{j})$ is the first eigenvalue on $\Omega_{j}$ with Dirichlet boundary conditions on $\partial\Omega_{j}$.
-If the $N$-uple of disjoint sets $(\Omega_{1},\dots,\Omega_{N})$ is a minimizer of (1.2) and if $u=(u_{1},\dots,u_{N})$ is the vector of the normalized first eigenfunctions (each $u_{j}$ is extended by zero outside $\Omega_{j}$), then the free interface
-$\mathcal{F}:=D\cap\Big{(}\cup_{j}\partial\Omega_{j}\Big{)}$
-can be decomposed as a regular part and a singular part as follows (see [CL07, Alp20]): the regular part $\mathrm{Reg}$ consists of all points $x\in\mathcal{F}$ of frequency $\gamma(u,x)=1$ and is a smooth manifold of dimension $d-1$; the singular part $\mathrm{Sing}$ is a closed $(d-2)$-rectifiable set composed of all the points of frequency $\gamma(u,x)>1$.
-We have the following theorem
-Theorem 1.7.
-Let $(\Omega_{1},\dots,\Omega_{N})$ be a minimizer of (1.2) and $u=(u_{1},\dots,u_{N})$ be the corresponding vector of the normalized first eigenfunctions.
-Then, the singular set of $\mathrm{Sing}$ of the free interface $\mathcal{F}$ can be decomposed as the disjoint union of
-$\mathrm{Sing}=\Big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
-The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}:=\big{\{}x\in\mathcal{F}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold.
-Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$, the set $\mathcal{F}$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}$.
-2. Optimal gap
+3. Singular blow-ups of minimal frequency
+The aim of this section is to classify homogeneous singular solutions with minimal frequency $\nicefrac{{3}}{{2}}$, i.e. functions in $\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$.
 3.1. A topological lemma
-Let $\mathcal{M}$ be a smooth $m$-dimensional manifold.
-We recall that a subset $\mathcal{F}\subset\mathcal{M}$ is a $C^{1}$ submanifold of dimension $m-1$ if for every $x_{0}\in\mathcal{F}$ there is an open neighborhood $\omega\subset\mathcal{M}$ of $x_{0}$ and a diffeomorphism $\Phi:\omega\to B_{1}\times(-1,1)$ such that $\Phi(\mathcal{F})$ is the graph $\Gamma_{\eta}$ of a $C^{1}$ function $\eta:B_{1}\to(-1,1)$; here $B_{1}$ is the unit ball in $\mathbb{R}^{m-1}$.
+This section is dedicated to a topological lemma, which we use both in the classification of the $\nicefrac{{3}}{{2}}$-homogeneous blow-ups (Proposition 3.5) and in the proof of the no-holes lemma (Lemma 6.6).
 Definition 3.1.
-Let $\mathcal{M}$ be a smooth $m$-dimensional manifold and let $\mathcal{F}\subset\mathcal{M}$ be a $C^{1}$ submanifold of dimension $m-1$.
-Given $x\in\mathcal{M}\setminus\mathcal{F}$ and $y\in\mathcal{M}\setminus\mathcal{F}$, we denote by $C(x\to y)$ the family of piecewise $C^{1}$ curves transversal to $\mathcal{F}$.
+Let $\mathcal{M}$ be a smooth $m$-dimensional manifold and let $\mathcal{F}\subset\mathcal{M}$ be a $C^{1}$ submanifold of dimension $m-1$ (that is, for every $x_{0}\in\mathcal{F}$ there is an open neighborhood $\omega\subset\mathcal{M}$ of $x_{0}$ and a $C^{1}$-diffeomorphism $\Phi:\omega\to B_{1}\subset\mathbb{R}^{d}$ such that $\Phi(\mathcal{F})=B_{1}\cap\{x_{m}=0\}$).
+For any $x,y\in\mathcal{M}\setminus\mathcal{F}$, we denote by $C(x\to y)$ the family of piecewise $C^{1}$ curves transversal to $\mathcal{F}$.
 Precisely, we say that $\ell\in\mathcal{C}(x\to y)$ if:
@@
   2. (Q2)
-    $Q_{i,j}$ is contained in an open set $\omega_{2}\subset\mathcal{M}$ for which there is a diffeomorphism $\Phi:\omega_{2}\to B_{1}\times(-1,1)$ such that $\Phi(\mathcal{F})$ is the graph $\Gamma_{\eta}$ of a $C^{1}$ function $\eta:B_{1}\to(-1,1)$, $B_{1}$ being the unit ball in $\mathbb{R}^{m-1}$.
+    $Q_{i,j}$ is contained in an open set $\omega_{2}\subset\mathcal{M}$ for which there is a $C^{1}$-diffeomorphism $\Phi:\omega_{2}\to B_{1}\subset\mathbb{R}^{m}$ such that $\Phi(\mathcal{F})=B_{1}\cap\{x_{m}=0\}$.
 We next notice that by perturbing slightly each of the points $X_{i,j}$ we can construct another family of $n\times n$ points
@@
 ∎
-2.2. Proof of Theorem 1.2
-In this section we prove Theorem 1.2. The proof is based on a dimension reduction argument and on the following
+3.2. Classification of the blow-ups of frequency $3/2$
+In this section we prove that any $\nicefrac{{3}}{{2}}$-homogeneous function $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$ is of the form $Y$ (see Proposition 3.5).
+We stress that even if this result is not explicitly stated in [ST15], the key elements (the dimension reduction argument and Lemma 3.4) are already contained in the proof of [ST15, Lemma 4.2]; here below we give the details since we need Proposition 3.5 in Theorem 1.1.
 Lemma 3.4.
-Let $d\geq 3$ and $N\geq 2$. Suppose that $u=(u_{1},\dots,u_{N})\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ is a non-trivial $\gamma$-homogeneous solution of the optimal partition problem, i.e.
-$u(x)=|x|^{\gamma}u\left(\frac{x}{|x|}\right)\quad\text{and}\quad\left\{\begin{aligned} -\Delta u_{i}&\leq 0,&&\text{in }\mathbb{R}^{d},\\ -\Delta(\sigma_{i}\cdot u)&\geq 0,&&\text{in }\mathbb{R}^{d},\end{aligned}\right.$
-in a distributional sense, for all $i=1,\dots,N$. Suppose that all the points on the unit sphere are regular, that is
+Let $d\geq 3$ and $N\geq 2$.
+Suppose that $u=(u_{1},\dots,u_{N})\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ is a non-trivial $\gamma$-homogeneous function.
+Suppose that all the free boundary points on the unit sphere are regular, that is
 $\mathcal{F}(u)\equiv\mathcal{F}_{1}(u)\quad\text{on}\quad\partial B_{1},$
@@
 $(\sigma\cdot u)(x):=\sum_{i=1}^{N}\sigma(i)u_{i}(x)$
-is harmonic in $\mathbb{R}^{d}$.
+is harmonic in $\mathbb{R}^{d}$. In particular, $\gamma$ is an integer.
 Proof.
@@
 In order to prove the harmonicity of $\sigma\cdot u$ we notice that if $B_{r}(x_{0})$ is a ball such that $\mathcal{F}(u)\cap B_{r}(x_{0})=\partial\Omega_{i}\cap\partial\Omega_{j}\cap B_{r}(x_{0})$ is a smooth graph in $B_{r}(x_{0})$, then for every pair of points $x_{i}\in\Omega_{i}\cap B_{r}(x_{0})$ and $x_{j}\in\Omega_{j}\cap B_{r}(x_{0})$, we can find a curve in $\mathcal{C}(x_{i}\to x_{j})$ that crosses $\mathcal{F}(u)$ only once.
-Thus, by construction
-$\sigma(x_{i})=-\sigma(x_{j})$
-and so the function $\sigma\cdot u$ is harmonic across $\mathcal{F}(u)$. This concludes the proof of Lemma 2.4. ∎
-We are now in position to prove our first main theorem.
-Proof of Theorem 1.2.
-In any dimension $d\geq 2$ and for any $N\geq 2$, we define:
-(2.2) $A(d,N):=\inf\Big{\{}\gamma>1\ :\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)\neq\emptyset\Big{\}}.$
-Since the admissible frequencies in dimension $d=2$ are classified, we have:
-$A(2,N)=\frac{3}{2}\qquad\text{and}\qquad 1<A(d,N)\leq\frac{3}{2}\quad\text{when}\quad d\geq 2.$
-We will show that $A(d,N)=3/2$ for every $d\geq 2$ by proving that
-(2.3) $A(d-1,N)\leq A(d,N)\quad\text{for all }d\geq 2.$
-In order to do this, we proceed in three steps.
-Step 1.
-The infimum in (2.2) is reached.
-Indeed, suppose that there is a sequence of $\gamma_{n}$-homogeneous global solutions $u^{n}\in\mathcal{S}_{\gamma_{n}}(\mathbb{R}^{d};N)$ such that $\gamma_{n}\to A(d,N)$.
-Up to a rescaling, we suppose that:
-$\|u^{n}\|_{L^{2}(\partial B_{1})}=1\qquad\text{and}\qquad\|\nabla u^{n}\|_{L^{2}(B_{1})}=\gamma_{n}.$
-Up to extracting a subsequence, we can suppose that $u^{n}$ converges strongly in $L^{2}(B_{1};\mathbb{R}^{N})$ and $L^{2}(\partial B_{1};\mathbb{R}^{N})$, and weakly in $H^{1}(B_{1};\mathbb{R}^{N})$ to some $A(d,N)$-homogeneous global solution $u\in\mathcal{S}_{A(d,N)}(\mathbb{R}^{d},N)$.
-Step 2.
-$\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}\neq\emptyset.$ Take any $k\in\{1,\dots,N\}$ and any connected component $\omega_{k}$ of $\Omega_{k}$.
-Since the component $u_{k}$ is homogeneous and non-negative in $\mathbb{R}^{d}$, its trace on $\omega_{k}\cap\partial B_{1}$ is (a multiple of) the first eigenfunction of the spherical Laplacian on $\omega_{k}\cap\partial B_{1}$ with eigenvalue
-$\lambda:=\gamma(\gamma+d-2)\quad\text{where}\quad\gamma=A(d,N),$
-with Dirichlet boundary conditions on $\partial\omega_{k}\cap\partial B_{1}$.
-In particular, by the Faber-Krahn inequality, the area of each $\omega_{k}\cap\partial B_{1}$ is bounded from below by a constant depending only on $\gamma$ and $d$, so the number of connected components of $\Omega_{k}$ is finite.
-Let $\widetilde{N}$ be the number of all connected components of $\mathbb{R}^{d}\setminus\mathcal{F}(u)$.
-We define the $\gamma$-homogeneous $\widetilde{u}\in\mathcal{A}(\mathbb{R}^{d};\widetilde{N})$ as the vector whose components are the restrictions of $u$ to the connected components of $\Omega_{k}$, $k=1,\dots,N$.
-It is immediate to check that $\widetilde{u}$ is still a global solution of the optimal partition problem, i.e.
-$\widetilde{u}\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};\widetilde{N})$.
-We now claim that $\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}\neq\emptyset$.
-If not, since by construction all $\widetilde{\Omega}_{j}:=\{\widetilde{u}_{j}>0\}$, $1\leq j\leq\widetilde{N}$ are connected, we can apply Lemma 2.4; thus, there is a map $\sigma:\{1,\dots,\widetilde{N}\}\to\{-1,1\}$ such that $\sigma\cdot\widetilde{u}$ is harmonic and $A(d,N)$-homogeneous in $\mathbb{R}^{d}$.
-Thus $A(d,N)$ is an integer and, since $A(d,N)>1$, we get that $A(d,N)\geq 2$.
-This is a contradiction with the bound $A(d,N)\leq 3/2$.
-Step 3.
-Dimension reduction.
-From the previous step we know that there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$.
-Without loss of generality, we can assume that $x_{0}=\vec{\bm{e}}_{1}$.
-By the definition of $A(d,N)$ we know that
-$A(d,N)\leq\gamma(u,\vec{\bm{e}}_{1}).$
-On the other hand, the homogeneity of $u$ implies that
-$\gamma(u,t\vec{\bm{e}}_{1})=\gamma(u,\vec{\bm{e}}_{1})\quad\text{for every}\quad t>0.$
-In particular, by the upper semicontinuity of $\gamma(u,\cdot)$ and the homogeneity of $u$, we have
-$A(d,N)=\gamma(u,0)\geq\lim_{t\to 0^{+}}\gamma(u,t\vec{\bm{e}}_{1})=\gamma(u,\vec{\bm{e}}_{1}).$
-Thus, we get
-$A(d,N)=\gamma(u,0)=\gamma(u,\vec{\bm{e}}_{1}),$
-but this implies that $u$ is invariant in the direction of $\vec{\bm{e}}_{1}$, that is:
-$u(x+t\vec{\bm{e}}_{1})=u(x)\quad\text{for every}\quad t\in\mathbb{R}.$
-But then the nontrivial function
-$\widetilde{u}(x_{2},\dots,x_{d}):=u(0,x_{2},\dots,x_{d}),$
-belongs to $\mathcal{S}_{A(d;N)}(\mathbb{R}^{d-1};N)$. This implies that $A(d-1;N)\leq A(d;N)$, thus proving (2.3) and so (2.2).
-In order to conclude the proof of Theorem 1.2, we are left to classify the $\nicefrac{{3}}{{2}}$-homogeneous global solutions.
+Thus, by construction $\sigma(x_{i})=-\sigma(x_{j})$ and so the function $\sigma\cdot u$ is harmonic across $\mathcal{F}(u)$.
+This concludes the proof of Lemma 3.4.
+∎
+Proposition 3.5.
 Let $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$.
-By Step 3 there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$.
-But then, reasoning as in Step 3, we have that
-$\gamma(u,0)=\gamma(u,x_{0})=\frac{3}{2}$
-and that $u$ is invariant in the direction of $x_{0}$.
+Then, up to relabeling the components of $u$, we have that $u_{j}\equiv 0$ for $j\geq 4$, and, up to a rotation of the coordinate system, $(u_{1},u_{2},u_{3})$ is of the form $(u_{1},u_{2},u_{3})=cY$ for some constant $c\in\mathbb{R}$ and $Y=(Y_{1},Y_{2},Y_{3})$ as in (1.3).
+Proof.
+Let $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$.
+By Lemma 3.4, there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$.
+But then, the upper semicontinuity of $\gamma(u,\cdot)$ and the homogeneity of $u$ give
+$\frac{3}{2}=\gamma(u,0)\geq\lim_{t\to+\infty}\gamma(u,tx_{0})=\gamma(u,x_{0}).$
+Since, by [ST15], $\nicefrac{{3}}{{2}}$ is the minimal frequency, we get that
+$\frac{3}{2}=\gamma(u,0)=\gamma(u,x_{0}),$
+which implies that $u$ is invariant in the direction of $x_{0}$.
 Up to a rotation of the coordinate system, we can suppose that $x_{0}=\vec{\bm{e}}_{1}$.
@@
 where $w$ is a $\nicefrac{{3}}{{2}}$-homogeneous solution in dimension 2. This gives the specific form of $w$ and the fact that $w$ has exactly three non-zero components. ∎
-3. Optimal partition of the sphere
-The purpose of this section is to prove Theorem 1.5, essentially basing ourselves on the results obtained in Section 2, namely Theorem 1.2 (and in [CTV05b, HHT09]).
-By [CTV05b, Theorem 3.1] we know that there exists $\phi\in H^{1}(\partial B_{1};\Sigma_{3})\cap C^{0,1}(B_{1};\mathbb{R}^{3})$ such that $\phi_{i}\not\equiv 0$ and $(\omega_{1},\omega_{2},\omega_{3})$, defined as $\omega_{i}:=\{\phi_{i}>0\}$, achieve
-(3.1) $\mathcal{L}_{3}:=\min\left\{\max_{i=1,2,3}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t.
-}\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\}.$
-Moreover, [CTV05b, Theorem 3.1] it is also proved that $\phi\in\mathcal{S}(\partial B_{1},3)$, in the sense that
-$-\Delta_{\partial B_{1}}\phi_{i}&\leq 0,&&\text{in }\partial B_{1},\\ -\Delta_{\partial B_{1}}(\sigma_{i}\cdot\phi)&\geq 0,&&\text{in }\partial B_{1}.\end{aligned}\right.$
-Now we recall [HHT09, Remark 3.14] which says that all the first eigenvalues are equal, namely $\lambda_{1}(\omega_{i})=\mathcal{L}_{3}$ for all $i=1,2,3$; we then let $\gamma>0$ be such that
-$\gamma(\gamma+d-2)=\mathcal{L}_{3}.$
-We now consider the $\gamma$-homogeneous extensions of $\phi_{i}$, that is $\varphi_{i}(r,\theta):=r^{\gamma}\phi_{i}(\theta)$.
-One can easily check that $\varphi\in H^{1}_{\textup{loc}}(\mathbb{R}^{d};\Sigma_{3})\cap C^{0,1}_{\textup{loc}}(\mathbb{R}^{d};\mathbb{R}^{3})$ belongs to the class $\mathcal{S}_{\gamma}(\mathbb{R}^{d},3)$, i.e.
-$-\Delta\varphi_{i}&\leq 0,&&\text{in }\mathbb{R}^{d},\\ -\Delta(\sigma_{i}\cdot\varphi)&\geq 0,&&\text{in }\mathbb{R}^{d}.\end{aligned}\right.$
-At this point, the proof of Theorem 1.5 follows easily. Indeed, one can test (3.1) with the $Y$-configuration and directly obtain that
-$\mathcal{L}_{3}\leq\frac{3}{2}\left(\frac{3}{2}+d-2\right),$
-thus implying that $\gamma\leq\nicefrac{{3}}{{2}}$.
-On the other hand, since $N=3$, by applying Theorem 1.2 to $\varphi$, we obtain that $\gamma=\nicefrac{{3}}{{2}}$ and that $\varphi$ coincides with the $Y$-configuration.
+We notice that as a consequence of Proposition 3.5, we obtain the uniqueness of the minimizer (in any dimension) to the optimal 3-partition of the sphere for the min-max problem studied in [CTV05a, HHT09, HHT10, ST15].
+Corollary 3.6.
+The $Y$-configuration is the unique minimizer of
+$\mathcal{L}_{3}:=\min\left\{\max_{i\in\{1,2,3\}}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t.
+}\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\},$
+where $\lambda_{1}(\omega_{i})$ is the first eigenvalue of the spherical Dirichlet Laplacian on $\omega_{i}$.
+Proof.
+Let $(\omega_{1},\omega_{2},\omega_{3})$ be a triple achieving the minimum (see for instance [HHT09]).
+We know that $\omega_{i}\subset\partial B_{1}$ is measurable and $\mathcal{H}^{d-1}(\omega_{i})>0$, for every $i$.
+Moreover, if $\phi_{i}$ is a first normalized eigenfunction on $\omega_{i}$, then by [HHT09, Theorem 3.4], there are constants $a_{i}\geq 0$, $i=1,2,3$, such that $\psi:=(a_{1}\phi_{1},a_{2}\phi_{2},a_{3}\phi_{3})$ satisfies
+$-\Delta\psi_{i}&\leq\mathcal{L}_{3}\psi_{i},&&\text{in }\partial B_{1},\\ -\Delta\Big{(}\psi_{i}-\sum_{j\neq i}\psi_{j}\Big{)}&\geq\mathcal{L}_{3}\Big{(}\psi_{i}-\sum_{j\neq i}\psi_{j}\Big{)},&&\text{in }\partial B_{1}.\end{aligned}\right.$
+By [HHT09, Theorem 3.4] that $(a_{1},a_{2},a_{3})\neq(0,0,0)$.
+We now claim that $a_{i}>0$ for all $i$.
+Suppose by contradiction that $a_{3}=0$, so that also $\psi_{3}=0$.
+By, the equation above, we get that $-\Delta(\psi_{1}-\psi_{2})=\mathcal{L}_{3}\,(\psi_{1}-\psi_{2})$ on $\partial B_{1}$.
+By the unique continuation principle for eigenfunctions, we get that $\mathcal{H}^{d-1}(\{\psi_{1}-\psi_{2}=0\})=0$, which leads to a contradiction since $\mathcal{H}^{d-1}(\omega_{3})>0$.
+Now, in view of [HHT09, Remark 3.14 (d)] and [ST15, Theorem 1.10], we know that $\mathcal{L}_{3}=\lambda_{1}(\omega_{i})=\frac{3}{2}\left(\frac{3}{2}+d-2\right)$ for all $i$.
+Thus, the $\nicefrac{{3}}{{2}}$ homogeneous extension $u$ of $\psi$ belongs to the class $\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$, so the conclusion follows from Proposition 3.5.
+∎
 4. Linearization around triple junctions
@@
 In this section we complete the proof of Theorem 1.1 by proving the regularity of the free interface $\mathcal{F}(u)$ around points of frequency $\nicefrac{{3}}{{2}}$.
-The main ingredients are the classification of the $3/2$-homogeneous blow-ups (Theorem 1.2) and the epiperimetric inequality Theorem 1.3.
+The main ingredients are the classification of the $\nicefrac{{3}}{{2}}$-homogeneous blow-ups (see Proposition 3.5) and the epiperimetric inequality Theorem 1.2.
 6.1. Rate of convergence of the blow-up sequence
@@
 In order to do so, we need to show that the conditions of Theorem 1.2 are fulfilled uniformly, at all points $x_{0}$ and at all small scales $r$, by multiples of $u^{r,x_{0}}$.
-We start with the following
+We start with the following.
 Lemma 6.1 (Uniform Hausdorff distance estimate).
@@
 Lemma 6.2 (Uniform convergence of rescalings with variable center).
-Let $u\in\mathcal{M}(D;N)$ be such that $\gamma(u,0)=\nicefrac{{3}}{{2}}$.
+Let $u\in\mathcal{M}(B_{1};N)$ be such that $\gamma(u,0)=\nicefrac{{3}}{{2}}$.
 Then, for all $\varepsilon>0$ there exists $\rho,r_{0}>0$ (depending on $d$, $N$ and $\varepsilon$) such that for all $z\in B_{\rho}\cap\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and for all $r<r_{0}$ there exists $Y\in\mathcal{M}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$ such that
-$\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{z,r},Y)\right\lVert_{L^{\infty}(B_{\nicefrac{{1}}{{2}}})}\leq\varepsilon,$
-where $u^{z,r}(x):=H(u,z,r)^{-\nicefrac{{1}}{{2}}}u(rx+z)$.
+$\left\lVert\mathrm{d}_{\Sigma_{N}}(u_{z,r},Y)\right\lVert_{L^{\infty}(B_{\nicefrac{{1}}{{2}}})}\leq\varepsilon,$
+where $u_{z,r}(x):=H(u,z,r)^{-\nicefrac{{1}}{{2}}}u(rx+z)$.
 Proof.
@@
 First of all, we observe that
-(6.4) $\sum_{i=1}^{N}\int_{B_{1}}|\nabla u^{z_{n},r_{n}}_{i}|^{2}\,\mathrm{d}x=\mathcal{N}(u^{z_{n},r_{n}},0,1).$
+(6.4) $\sum_{i=1}^{N}\int_{B_{1}}|\nabla(u_{z_{n},r_{n}})_{i}|^{2}\,\mathrm{d}x=\mathcal{N}(u_{z_{n},r_{n}},0,1).$
 Thus, up to a subsequence
-(6.5) $u^{z_{n},r_{n}}\to u_{0}\quad\text{strongly in }H^{1}(B_{1};\Sigma_{N})~{}\text{and uniformly in }B_{\nicefrac{{1}}{{2}}},$
-for some $u_{0}\in H^{1}(B_{1};\Sigma_{N})$.
-Moreover, since $u^{z_{n},r_{n}}\in\mathcal{M}(B_{1};N)$ for all $n$, also $u_{0}\in\mathcal{M}(B_{1};N)$.
+(6.5) $u_{z_{n},r_{n}}\to u_{0}\quad\text{strongly in }H^{1}(B_{1};\Sigma_{N})~{}\text{and uniformly in }B_{\nicefrac{{1}}{{2}}},$
+for some $u_{0}\in H^{1}(B_{1};\Sigma_{N})$, $u_{0}\not\equiv 0$.
+Moreover, since $u_{z_{n},r_{n}}\in\mathcal{M}(B_{1};N)$ for all $n$, also $u_{0}\in\mathcal{M}(B_{1};N)$.
 We now observe that, since $z_{n}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and by monotonicity, we have
-$\frac{3}{2}=\mathcal{N}(u,z_{n},0^{+})\leq\mathcal{N}(u,z_{n},Rr_{n})=\mathcal{N}(u^{z_{n},r_{n}},0,R),$
+$\frac{3}{2}=\mathcal{N}(u,z_{n},0^{+})\leq\mathcal{N}(u,z_{n},Rr_{n})=\mathcal{N}(u_{z_{n},r_{n}},0,R),$
 for any $R>0$ and $n$ sufficiently large. Passing to the limit as $n\to\infty$, we get
 $\frac{3}{2}\leq\mathcal{N}(u_{0},0,R)\quad\text{for all }R>0.$
-By (6.5) and scaling properties, we have that
-$\mathcal{N}(u_{0},0,R)=\lim_{n\to\infty}\mathcal{N}(u^{z_{n},r_{n}},0,R)=\lim_{n\to\infty}\mathcal{N}(u,z_{n},Rr_{n}),$
-for all $R>0$. Therefore, by upper semicontinuity and monotonicity of the frequency function and Lemma 2.4, we have that for all $R,\tau>0$ sufficiently small, there holds
+By the continuity of $x\mapsto\mathcal{N}(u,x,r)$, the monotonicity of $r\mapsto\mathcal{N}(u,x,r)$ and Lemma 3.4, we have that for all $R,\tau>0$ sufficiently small, there holds
 $\frac{3}{2}\leq\mathcal{N}(u_{0},0,R)\leq\mathcal{N}(u,0,\tau R).$
@@
 Proposition 6.3.
-Let $\alpha:=\varepsilon(d+1)$, with $\varepsilon>0$ being as in Theorem 1.3.
-Let $D\subseteq\mathbb{R}^{d}$ be open and let $u\in\mathcal{M}(D;N)$.
-Then, for all compact $K\subseteq D$ there exists $R>0$ such that for all $x_{0}\in K$ satisfying $\gamma(u,x_{0})=\nicefrac{{3}}{{2}}$, there exists $C_{\textup{rate}}>0$ depending on $d$ and $\mathcal{N}(u,x_{0},R)$ such that
+Let $\alpha:=\varepsilon(d+1)$, with $\varepsilon>0$ being as in Theorem 1.2 and let $u\in\mathcal{M}(B_{1};N)$.
+Then, for all compact $K\subseteq B_{1}$ there exists $R>0$ such that for all $x_{0}\in K\cap\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, there exists $C_{\textup{rate}}>0$ depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,\mathrm{d}_{\mathcal{H}}(K,\partial B_{1}))$ such that
 $\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{x_{0},r_{2}},u^{x_{0},r_{1}})\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq\sum_{i=1}^{N}\left\lVert u_{i}^{x_{0},r_{2}}-u_{i}^{x_{0},r_{1}}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq C_{\textup{rate}}(r_{2}^{\alpha}-r_{1}^{\alpha})$
@@
 $\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{x_{0},r},Y^{x_{0}})\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq\sum_{i=1}^{N}\left\lVert u_{i}^{x_{0},r}-Y^{x_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq C_{\textup{rate}}r^{\alpha},$
-for all $r<\operatorname{\mathrm{dist}}(x_{0},\partial D)$.
+for all $r<R$.
 Proof.
@@
 ∎
-As a corollary, we also obtain the following
+As a corollary, we also obtain the following.
 Proposition 6.4 (Oscillation of the $\nicefrac{{3}}{{2}}$-blow-up limits).
-Let $\bar{\alpha}:=2\alpha/(\alpha+3)$, where $\alpha$ is as in Proposition 6.3 and let $u\in\mathcal{M}(D;N)$.
-Then, for all compact $K\subseteq D$ there exists $r_{0}>0$ (depending on $d$ and $K$) and $C_{\textup{osc}}>0$ (depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$) such that
+Let $\bar{\alpha}:=2\alpha/(\alpha+3)$, where $\alpha$ is as in Proposition 6.3 and let $u\in\mathcal{M}(B_{1};N)$.
+Then, for all compact $K\subseteq B_{1}$ there exists $r_{0}>0$ (depending on $d$ and $K$) and $C_{\textup{osc}}>0$ (depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,\mathrm{d}_{\mathcal{H}}(K,\partial B_{1}))$) such that
 $\sum_{i=1}^{3}\left\lVert Y^{x_{0}}_{i}-Y^{z_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\leq C_{\textup{osc}}|x_{0}-z_{0}|^{\bar{\alpha}}$
@@
 (6.7) $\sum_{i=1}^{N}\left(\left\lVert u^{x_{0},r}_{i}-Y^{x_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}+\left\lVert u^{z_{0},r}_{i}-Y^{z_{0}}_{i}\right\lVert_{L^{2}(\partial B_{1})}^{2}\right)\leq C\,r^{\alpha},$
-for $r$ sufficiently small depending on $d$ and $\mathrm{d}_{\mathcal{H}}(K,\partial D)$ and $C>0$ depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$.
+for $r$ sufficiently small depending on $d$ and $\mathrm{d}_{\mathcal{H}}(K,\partial B_{1})$ and $C>0$ depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$.
 On the other hand, if $L_{K}>0$ denotes the Lipschitz constant of $u$ in $K$, we have that
@@
 for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$, we derive that $\gamma(u_{n},x)<\nicefrac{{3}}{{2}}$ for $n$ sufficiently large for all $x\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$.
-On the other hand, from Theorem 1.2 we obtain that $\gamma(u_{n},x)=1$ for all $x=(x^{\prime\prime}_{0},x_{d-1},x_{d})\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$.
+On the other hand, by the optimal frequency gap we also know that $\gamma(u_{n},x)=1$ for all $x=(x^{\prime\prime}_{0},x_{d-1},x_{d})\in B_{\rho}\cap C_{\varepsilon}\cap\mathcal{F}(u_{n})$.
 As a consequence, for some $n$ large enough, we can find $t>0$ such that in the set
@@
 Proof of Theorem 1.1.
-Let $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}$ and let $u_{0}$ be the $\nicefrac{{3}}{{2}}$-homogeneous blow-up of $u$ at $x_{0}$, which is unique and non-zero by Proposition 6.3.
+Let $x_{0}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ and let $u_{0}$ be the $\nicefrac{{3}}{{2}}$-homogeneous blow-up of $u$ at $x_{0}$, which is unique and non-zero by Proposition 6.3.
 Up to multiplying $u$ with a constant, we can suppose that $u_{0}=Y$.
-By Lemma 6.5 and Theorem 1.2 we know that in a neighborhood of $x_{0}$ there are only points of frequency $1$ or $\nicefrac{{3}}{{2}}$.
-By the no-holes lemma Lemma 6.6, we know that for every $x^{\prime\prime}\in\mathbb{R}^{d-2}$ with $|x^{\prime\prime}|$ small enough, there exist $(x_{d-1},x_{d})\in\mathbb{R}^{2}$ such that $x=(x^{\prime\prime},x_{d-1},x_{d})\in\nicefrac{{3}}{{2}}$.
-By Proposition 6.3, we know that the point $(x_{d-1},x_{d})$ is unique, so in a neighborhood of $x_{0}$, $\mathcal{F}_{\nicefrac{{3}}{{2}}}$ is a graph of a function $\eta:\mathbb{R}^{d-2}\to\mathbb{R}^{2}$.
+By Lemma 6.5 we know that in a neighborhood of $x_{0}$ there are only points of frequency $1$ or $\nicefrac{{3}}{{2}}$.
+By the no-holes lemma Lemma 6.6, we know that for every $x^{\prime\prime}\in\mathbb{R}^{d-2}$ with $|x^{\prime\prime}|$ small enough, there exist $(x_{d-1},x_{d})\in\mathbb{R}^{2}$ such that $x=(x^{\prime\prime},x_{d-1},x_{d})\in{\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)}$.
+By Proposition 6.3, we know that the point $(x_{d-1},x_{d})$ is unique, so in a neighborhood of $x_{0}$, $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ is a graph of a function $\eta:\mathbb{R}^{d-2}\to\mathbb{R}^{2}$.
 The $C^{1,\alpha}$-regularity of $\eta$ follows from Proposition 6.3, Proposition 6.4 and Lemma 6.1 by a standard argument.
-Finally, the regularity of $\mathcal{F}(u)$ around $x_{0}$ is a consequence of the regularity of $\eta$, the uniqueness of the blow-up and the epsilon-regularity for $\mathcal{F}_{1}$ ([OV24, Lemma 9.8]); the argument is standard and for the details we refer for instance to [OV24, Section 9].
+Finally, the regularity of $\mathcal{F}(u)$ around $x_{0}$ is a consequence of the regularity of $\eta$, the uniqueness of the blow-up and the epsilon-regularity for $\mathcal{F}_{1}{(u)}$ ([OV24, Lemma 9.8]); the argument is standard and for the details we refer for instance to [OV24, Section 9].
 ∎

### v3

@@
 $\mathrm{d}_{\Sigma_{N}}(X,Y):=\begin{cases}|X_{i}-Y_{i}|,&\text{if }X_{j}=Y_{j}=0~{}\text{for all }j\neq i,\\ |X_{i}|+|Y_{j}|,&\text{if }X_{i},Y_{j}>0~{}\text{for some }i\neq j.\end{cases}$
-By the works of Gromov-Schoen [GS92], Caffarelli-Lin [CL07, CL08], Tavares-Terracini [TT12] and Soave-Terracini [ST15] it is known that if
+By the works of Gromov-Schoen [GS92], Sun [Sun03], Caffarelli-Lin [CL07, CL08], Tavares-Terracini [TT12] and Soave-Terracini [ST15] it is known that if
 $\mathcal{F}(u)=\{x\in B_{1}\ :\ u(x)=0\},$
@@
 $\mathrm{Reg}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=1\big{\}}\qquad\text{and}\qquad\mathrm{Sing}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)\geq\nicefrac{{3}}{{2}}\big{\}}.$
-By [GS92, CL08], in every dimension $d\geq 2$, the regular part $\mathrm{Reg}(u)$ is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92]) and is $(d-2)$-rectifiable ([Alp20, Dee22]).
+By [GS92, Sun03, CL08], in every dimension $d\geq 2$, the regular part $\mathrm{Reg}(u)$ is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92, Sun03] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92, Sun03]) and is $(d-2)$-rectifiable ([Alp20, Dee22]).
 In dimension $d=2$, Conti, Terracini and Verzini showed in [CTV03] that the singular set is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point (see also the recent paper [LMNS24] for the case of generic solutions).
@@
 Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, the nodal set $\mathcal{F}(u)$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$.
-Results in the spirit of Theorem 1.1 exist in the context of area-minimizing currents, where the nodal set $\mathcal{F}(u)$ is replaced by the support of an area-minimizing current (mod $p=3$); in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93], in any dimension $d\geq 3$ (see also the lecture notes [Min24]); we also remark that the analysis for general co-dimension 1 area-minimizing currents mod p was completed only recently in [DHMS20], [DHM+21] and [DHM+22].
-To our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$.
+Results in the spirit of Theorem 1.1 exist in the context of area-minimizing surfaces; in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93] in any dimension $d\geq 3$.
+The techniques from [Tay73, Tay76, Sim93] have been abundantly explored in the last decade (see for instance [DHMS20, DHM+21, DHM+22] and the references therein), but fine results as Theorem 1.1 have been obtained only for the area functional and its generalizations so far.
+In fact, to our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$.
 The main ingredient in the proof of Theorem 1.1 is the following epiperimetric inequality (Theorem 1.2) for the $\nicefrac{{3}}{{2}}$-Weiss energy
@@
 There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on the dimension $d$ such that the following holds.
-For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that
+For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $c_{i}\geq 0$ for all $i=1,\dots,N$,
 $\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{Y_{i}>0\}\right)+\sum_{i=4}^{N}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{x_{d-1}=x_{d}=0\}\right)\leq\tau$
@@
 $\left\lVert\mathrm{d}_{\Sigma_{N}}(c,Y)\right\lVert_{H^{1}(B_{1})}\leq\delta,$
-there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u=c$ on $\partial B_{1}$ and
+there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u_{i}\geq 0$ for all $i$, $u=c$ on $\partial B_{1}$, and
 $W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
-The proof of Theorem 1.2 is obtained via a contradiction argument in the spirit of Weiss [Wei99] (see also [FS16, GPS16]).
+The proof of Theorem 1.2 is obtained via a contradiction argument in the spirit of Weiss [Wei99] (see also [FS16, GPS16, Tay76, Tay73]).
 The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, new in this field.
+One of the main consequences of Theorem 1.2 is the uniqueness of the ($\nicefrac{{3}}{{2}}$-homogeneous) blow-up at every point of frequency $\nicefrac{{3}}{{2}}$.
+We notice that the uniqueness of the tangent maps was claimed in [CL10, Section 3].
+The argument in [CL10] is based on a Monneau-type formula obtained via the Weiss-energy linearization identity from [CL10, Lemma 3.2].
+Unfortunately, there is a missing term in [CL10, Lemma 3.2], which compromises the conclusion in [CL10].
+The correct formula is given in Lemma 5.3 and contains an extra term $\beta$.
+This term is linear and positive, and requires particular attention in the proof of the epiperimetric inequality.
+Finally, we also point out that the term $\beta$ from Lemma 5.3 is specific for our functional and contains information on the interaction between the different phases.
+Our epiperimetric inequality shares some common features with the epiperimetric inequalities from [Tay76, Tay73], for instance, the concept of “linearization” and the difficulties related to the “parametrization” over the limit cone.
+On the other hand, our functional is of different nature, which is reflected, in particular, in the classifications of the solutions of the linearized problem in Section 4, in the Weiss-energy linearization identity from Lemma 5.3, and in the constructions of the competitors in Step 1 and Step 4 of the proof of the epiperimetric inequality in Section 5, where we face structural problems caused by the segregation constraint which is not present in [Wei99, Tay76, Tay73].
+Once we have the epiperimetric inequality Theorem 1.2, the conclusion in Theorem 1.1 follows mainly from the classification of the $\nicefrac{{3}}{{2}}$-homogeneous tangent maps (Proposition 3.5), from the frequency gap from above (Lemma 6.5), and from a topological argument (Lemma 6.6) similar to [Sim93, Tay76, Tay73], which guarantees the persistence of singularities over the triple junction model.
 1.1. Two consequences of Theorem 1.1
@@
 Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, the set $\mathcal{F}{(u)}$ consists of three $(d-1)$-manifolds with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$.
+1.2. The Bishop-Friedland-Hayman’s conjecture for the min-max
+Consider, for $p\in(0,+\infty]$, the family of optimal partition problems
+($\text{OP}_{p}$) $\min\left\{\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t.
+}\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\},$
+where $\lambda_{1}(\omega)$ denotes the first Dirichlet eigenvalue of the spherical Laplacian in $\omega\subseteq\partial B_{1}$, and
+$\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}:=\max_{i\in\{1,\dots,N\}}\lambda_{1}(\omega_{i})\quad\text{when}\quad p=+\infty.$
+In the particular case of the $3$-partition, i.e. $N=3$, we have the following conjecture.
+Problem 1.5 (Bishop-Friedland-Hayman-Helffer-Hoffmann-Ostenhof-Terracini).
+For $N=3$, prove that the $Y$-configuration is the unique minimizer of ($\text{OP}_{p}$), for any $p\in(0,+\infty]$.
+This conjecture was posed in [Bis92] in the case $p=1$, but can be generalized for any $p\in(0,+\infty]$ (see [FH76] and [HHT10]) and is related to shape optimization and other optimal partition problems [CTV03, CTV05a, BBH98, Hel10, HH10, HHT10].
+1.5 was solved for $p=+\infty$ in dimension $d=3$ by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10].
+In Section 3.2 we solve 1.5, for $p=+\infty$ and in any dimension $d\geq 3$, as a consequence of the classification of $\nicefrac{{3}}{{2}}$-homogeneous tangent maps (first obtained in [ST15]), see Corollary 3.6.
+Organization of the paper
 The rest of the paper is organized as follows.
@@
 5. Epiperimetric inequality at points of frequency $\nicefrac{{3}}{{2}}$
+In this section we prove Theorem 1.2; we start with the following general lemma, which will be useful in the construction of the competitors in Step 1 and Step 4 of the proof.
 Lemma 5.1 (Distance on $\Sigma_{N}$).
@@
 If $u_{k}(x)=0$ for all $k=1,\dots,N$ the statement is trivial. ∎
+The following lemma contains a general integration by parts identity for tangent maps.
 Lemma 5.2.
@@
 where the constant $C$ depends on $d$, $\|v\|_{L^{\infty}}$, $\|\nabla v\|_{L^{\infty}}$, $\|u\|_{L^{\infty}}$, and $\|\nabla u\|_{L^{\infty}}$. ∎
-Lemma 5.3.
-For any $u\in H^{1}(B_{1};\Sigma_{N})$ and any $\gamma\geq 0$, we let
+The next lemma is the Weiss-energy linearization identity which is a key element in the argument in the proof of Theorem 1.2.
+Lemma 5.3 (Weiss-energy linearization identity).
+For any $u\in H^{1}(B_{1};\Sigma_{N})$ and $\gamma\geq 0$, let
 $W_{\gamma}(u):=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x-\gamma\sum_{i=1}^{N}\int_{\partial B_{1}}u_{i}^{2}\,\mathrm{d}S.$
@@
 thus concluding the proof. ∎
-Theorem 5.4.
-There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on $d$ such that the following holds.
-For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $c_{n,i}\geq 0$ for all $i=1,\dots,N$,
-$\sum_{i=1}^{3}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{Y_{i}>0\}\right)+\sum_{i=4}^{N}\mathrm{d}_{\mathcal{H}}\left(\{c_{i}>0\},\{x_{d-1}=x_{d}=0\}\right)\leq\tau$
-and
-$\left\lVert\mathrm{d}_{\Sigma_{N}}(c,Y)\right\lVert_{H^{1}(B_{1})}\leq\delta,$
-there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u_{i}\geq 0$ for all $i$, $u=c$ on $\partial B_{1}$, and
-$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
-Proof.
+We conclude this section with the following.
+Proof of Theorem 1.2.
 In the whole proof, with an abuse of notation, for the sake of simplicity, we write $\Omega_{i}$ in place of $\Omega_{i}\cap B_{1}$, for $i=1,2,3$.
@@
 acknowledges support from the projects PRA_2022_14 GeoDom (PRA 2022 - Università di Pisa) and MUR-PRIN “NO3” (n.2022R537CS).
+Corresponding author
+Bozhidar Velichkov (e-mail: bozhidar.velichkov@unipi.it)
+Conflict of interests and data availability statements
+The authors have no competing interests to declare. Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.